### v3

# Itô’s formula for flows of conditional measures on semimartingales

Abstract.
Motivated by recent developments of mean-field systems with common noise, this paper establishes Itô’s formula for flows of conditional probability measures under a common filtration associated with general semimartingales. It generalizes existing works on flows of conditional measures on Itô processes and flows of deterministic measures on general semimartingales. It is derived via the cylindrical function approach. One key technical component involves constructing conditionally independent copies of stochastic processes, which allows for establishing the equivalence between stochastic integrals with respect to the conditional laws of semimartingales and the conditional expectation of stochastic integrals with respect to copies of semimartingales.
As an application, a class of control problems for jump-diffusions with partial observations is analyzed.
1. Introduction
Itô’s formula
The classical Itô’s formula for semimartingales, the stochastic equivalent of the chain rule in calculus, is a cornerstone of stochastic analysis [36]. It intrinsically links partial differential equations (PDEs) with diffusion processes. Recent advancements in mean-field game and control theories [9, 11, 14] have led to its extension to flows of probability measures, initially for diffusions [9] and later for discontinuous semimartingales ([31, 53]). This generalized Itô’s formula is crucial for deriving master equations in mean-field games and Bellman dynamic programming equations for McKean–Vlasov control problems, see [14] and the references therein.
Most recent studies in mean-field theory involve controls and games with common noise (see for instance [1, 2, 5, 7, 12, 13, 17, 19, 24, 30, 34, 40, 41, 44, 45, 48, 54]) or stochastic jump processes ([10, 28, 32, 33, 35, 49]). In either case, Itô’s formula has been developed respectively in ([9, 14]) and ([31, 53]). In order to analyze rigorously the McKean–Vlasov dynamics with both semimartingales and idiosyncratic and common noises, it is necessary to establish the Itô’s formula for flows of conditional laws with semimartingales, which does not seem to exist, to the best of our knowledge. This is the primary focus of this paper.
Our work
[…]
One key step is to analyze the Itô integral $\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}^{\mathcal{G}}$ with $\{X_{t}^{\mathcal{G}}:={\mathbb{E}}[X_{t}|{\mathcal{G}}]\}_{t\in[0,T]}$ over some appropriate adapted process $\{\eta_{t}\}_{t\in[0,T]}$. This is crucial as the integral of the conditional expectation such as $\int_{0}^{t}\eta_{s}{\mathrm{d}}X^{\mathcal{G}}_{s}$ is not necessarily equivalent to the conditional expectation of the integral of the form ${\mathbb{E}}\big{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}|{\mathcal{G}}\big{]}$ (See discussions and examples in Section 3). Our analysis is built on two key technical components: one is to construct conditionally independent copies of general stochastic processes (Theorem 3.1), the other is to establish the equivalence between the conditional expectation of stochastic integrals with respect to semimartingales and stochastic integrals with respect to the conditional law of semimartingales, as well as the equivalence between the quadratic variation of the conditional expectation of a semimartingale and the conditional expectation of the quadratic variation of its two conditionally independent copies (Theorem 3.5).
With these key technical components, Itô’s formula for flows of conditional laws on semimartingales is established (Theorem 2.11): the result is first established for cylindrical functions, defined as smooth mean-field functions with integrable representations (see Definition 4.1). Since these functions are dense in the desired function space, the general form is then obtained by a localization argument combined with suitable versions of the dominated convergence theorem.
This Itô’s formula (Theorem 2.11) is applied to analyze control problems with partially observed jump-diffusions. An appropriate form of the verification theorem 6.2 is derived, which recovers the results in [4] as special cases.
Related works
One of the key technical elements is the construction of conditionally independent copies of stochastic processes. This concept has been treated informally, for instance, when defining a semimartingale by stochastic differential equations involving both idiosyncratic and common noise in [14].
Our construction of conditionally independent copies of general stochastic processes is inspired by the proofs of Yamada-Watanabe theorem (see for instance [39, 42]) and the theory of weak and strong solutions of stochastic differential equations (see for instance [37, 43]). Most existing methods specify the common information $\mathcal{G}$} as a stochastic input, and make use of the existence of the regular conditional distribution to construct an extended probability space. In contrast, we construct conditionally independent copies directly, rather than utilizing the regular conditional distribution. (See the discussion at the beginning of Section 5 for a detailed comparison).
Our construction approach generalizes the construction of conditionally independent copies to a broad class of semimartingales and allows for an arbitrary sub $\sigma$-algebra. This result may be of independent theoretical interest, besides its essential role in our derivation of Itô’s formula for flow of conditional laws on semimartingales.
A crucial assumption in our analysis is a condition on the filtration known as the compatibility condition. This condition appears in various contexts under different names, such as immersion [38], very good extensions [37], and natural extensions [50]. In the theory of filtration enlargement, it is known as the H-hypothesis ([8, 26]), which plays a central role for characterizing semimartingale properties with filtration expansion or shrinkage. Following [43], we adopt the term compatibility to be aligned with recent literature on mean-field games ([6, 14, 15, 16, 46]).
Finally, it is worth pointing out that there are several methods to derive Itô’s formula and its variants, including the time discretization approach in [9] and [49] for mean-field jump diffusions, the density approach in [11] using the Fokker–Planck equation, and the particle approximation approach of [14, 18], [25, 52] by flows of empirical measures. The cylindrical function approach has proved appropriate and powerful for analysis of general semimartingales. It is initially explored for Fleming–Viot processes [27], later adopted in the analysis of polynomial diffusions [23], and for Itô’s formulas on flow of measures with (discontinuous) semimartingales [31] and for non-anticipative maps of càdlàg rough paths [22].
Notation
Throughout the paper, we will adopt the following notations, unless otherwise specified.
  • We use the bold letter $\bm{x},\bm{X}$ to denote a (random) vector with multiple dimensions or a matrix depending on the context. $|\bm{X}|$ denotes the Euclidean norm of a vector $\bm{X}\in{\mathbb{R}}^{d}$, or the Frobenius norm of a matrix.
  • ${\mathcal{P}}({\mathbb{R}}^{d})$ denotes the space of probability measures on ${\mathbb{R}}^{d}$ with the topology of weak convergence. ${\mathcal{P}}_{p}({\mathbb{R}}^{d})$ for $p\in[1,\infty)$ denotes the probability measures with finite $p$-th moment, equipped with the Wasserstein-$p$ metric.
  • $C^{k}({\mathbb{R}}^{d})$ is the set of all $k$-th differentiable functions on ${\mathbb{R}}^{d}$ with continuous derivatives up to the $k$-th order, and $C_{b}^{k}({\mathbb{R}}^{d})\subset C^{k}({\mathbb{R}}^{d})$ is the set of all $k$-th differentiable functions on ${\mathbb{R}}^{d}$ with bounded and continuous derivatives up to the $k$-th order, with the convention $C({\mathbb{R}}^{d})=C^{0}({\mathbb{R}}^{d})$ and $C_{b}({\mathbb{R}}^{d})=C^{0}_{b}({\mathbb{R}}^{d})$.
  • For a $d$-dimensional vector-valued random variable $\bm{X}$, $\|\bm{X}\|_{L^{p}}$ represents its $p$-th moment, i.e., $\|\bm{X}\|_{L^{p}}={\mathbb{E}}[|\bm{X}|^{p}]^{\frac{1}{p}}$.
  • For vectors $\bm{a}=(a_{i})_{i=1,..,d},\bm{b}=(b_{i})_{i=1,..,d}\in{\mathbb{R}}^{d}$, denote $\bm{a}\cdot\bm{b}=\sum_{i=1}^{d}a_{i}b_{i}$. For matrices $\bm{C}=(C_{ij})_{i,j=1,..,d},\bm{D}=(D_{ij})_{i,j=1,..,d}\in{\mathbb{R}}^{d\times d}$, denote $\bm{C}:\bm{D}=\sum_{i,j=1}^{d}C_{ij}D_{ij}$.
  • For $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$ and $\varphi:{\mathbb{R}}^{d}\to{\mathbb{R}}$ such that $\int_{{\mathbb{R}}^{d}}|\varphi(\bm{x})|{\mathrm{d}}\mu(\bm{x})<\infty$, set
    $\langle\mu,\varphi\rangle:=\int_{{\mathbb{R}}^{d}}\varphi(\bm{x}){\mathrm{d}}\mu(\bm{x}).$
  • For a càdlàg process $\bm{X}_{t}$, we adopt the standard notation:
    $\bm{X}_{t-}:=\lim_{s\nearrow t}\bm{X}_{s},\quad\Delta\bm{X}_{t}:=\bm{X}_{t}-\bm{X}_{t-},$
    and $\bm{X}^{c}$ to represent the continuous part of the process.
2. Assumptions and main result
Fix $T>0$ and $d\in\mathbb{N}^{+}$. Consider two ${\mathbb{R}}^{d}$-valued $\mathbb{F}$}-semimartingales $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ and $\bm{Y}=(\bm{Y}_{t})_{t\in[0,T]}$ on a completed filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$, and $\mu=\big{\{}\mu_{t}:=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})\big{\}}_{t\in[0,T]}$ for a sub-filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$. Our focus is to establish an Itô’s formula for the functional $\Phi(\mu_{t},\bm{Y}_{t})$, with $\Phi:{\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ for some $p\in[2,\infty)$ under suitable regularity conditions.
2.1. Assumptions and preliminary tools
Throughout the paper, we will make the following assumptions on the semimartingales $\bm{X}$ and $\bm{Y}$, the filtrations $\mathbb{G}$} and $\mathbb{F}$}, and the function $\Phi$. These are standard assumptions to ensure that stochastic integrals are properly defined ([51]) with the presence of common noise ([47]), and for flow of measures on semimartingales ([21, 31]).
First, we specify the assumptions on the semimartingales $\bm{X}$ and $\bm{Y}$.
Assumption 2.1.
For some $1\leq p\leq\infty$,
  • $\big{\|}\bm{X}\big{\|}_{{\mathcal{H}}^{p}}=\inf_{\bm{M},\bm{V}}\bigg{\|}|\bm{X}_{0}|+\sqrt{[\bm{M},\bm{M}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}\bm{V}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty,$ (2.1)
    $\big{\|}\bm{Y}\big{\|}_{{\mathcal{H}}^{p}}=\inf_{\bm{N},\bm{U}}\bigg{\|}|\bm{Y}_{0}|+\sqrt{[\bm{N},\bm{N}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}\bm{U}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty,$
    where the infimum is taken over all possible decompositions
    $\bm{X}_{t}=\bm{X}_{0}+\bm{M}_{t}+\bm{V}_{t}\text{ and }\bm{Y}_{t}=\bm{Y}_{0}+\bm{N}_{t}+\bm{U}_{t},$
    where $\bm{V}=(\bm{V}_{t})_{t\in[0,T]}$ and $\bm{U}=(\bm{U}_{t})_{t\in[0,T]}$ are adapted càdlàg processes of finite variation with $\bm{V}_{0}=\bm{U}_{0}=\bm{0}$, and $\bm{M}=(\bm{M}_{t})_{t\in[0,T]}$ and $\bm{N}=(\bm{N}_{t})_{t\in[0,T]}$ are $\mathbb{F}$}-local martingales such that $\bm{M}_{0}=\bm{N}_{0}=\bm{0}$.
  • ${\mathbb{E}}\bigg{[}\Big{(}\sum_{0\leq s\leq T}|\Delta\bm{X}_{s}|\Big{)}^{p}\bigg{]}+{\mathbb{E}}\bigg{[}\Big{(}\sum_{0\leq s\leq T}|\Delta\bm{Y}_{s}|\Big{)}^{p}\bigg{]}<\infty.$ (2.2)
There are two propositions that are necessary for the subsequent analysis, especially for the localization argument in establishing the Itô’s lemma for the flow of conditional measures on semimartingales, as will be clear in Section 4.2.
Proposition 2.2 ([51, Section V, Theorem 2]).
For any $1\leq p\leq\infty$, define a norm on the space $D$ of all $\mathbb{R}$}-valued adapted càdlàg processes, such that for $H=(H_{t})_{t\in[0,T]}\in D$,
$\|H\|_{{\mathcal{S}}^{p}}:=\Big{\|}\sup_{0\leq t\leq T}|H_{t}|\Big{\|}_{L^{p}}.$
Then, there exists a constant $c_{p}$ depending only on $p$ such that
$\|H\|_{{\mathcal{S}}^{p}}\leq c_{p}\|H\|_{{\mathcal{H}}^{p}},$
for any $\mathbb{R}$}-valued semimartingale $H$.
Proposition 2.3 (Emery inequality, [51, Section V, Theorem 3]).
Let $Z$ be an $\mathbb{R}$}-valued semimartingale, $H$ be an $\mathbb{R}$}-valued adapted càdlàg process, and $\frac{1}{p}+\frac{1}{q}=\frac{1}{r}$ ($1\leq p,q\leq\infty)$, then
$\Big{\|}\int_{0}^{\cdot}H_{s-}{\mathrm{d}}Z_{s}\Big{\|}_{{\mathcal{H}}^{r}}\leq\|H\|_{{\mathcal{S}}^{p}}\|Z\|_{{\mathcal{H}}^{q}}.$
Remark 2.4.
Propositions 2.2 and 2.3 also hold for ${\mathbb{R}}^{d}$-valued semimartingales with appropriate forms in $d\in{\mathbb{N}}^{+}$.
Moreover, a proper form of dominated convergence theorem has been established as follows.
Lemma 2.5.
([51, Page 273, Lemma]) Let $p,q,r$ be given such that $\frac{1}{p}+\frac{1}{q}=\frac{1}{r}$ where $1<r<+\infty$. Let $Z$ be a semimartingale such that $\|Z\|_{\mathcal{H}_{q}}<+\infty$, and let $\{H^{n}=(H^{n}_{t})_{t\in[0,T]}\}_{n\in{\mathbb{N}}}$ and $Y=(Y_{t})_{t\in[0,T]}$ be semimartingales such that $\|H^{n}\|_{{\mathcal{S}}^{p}}<+\infty$ for all $n\in{\mathbb{N}}$, $\|Y\|_{{\mathcal{S}}^{p}}<+\infty$ and $|H^{n}_{t}|\leq Y_{t}$. Suppose $\lim_{n\to\infty}H^{n}_{t-}(\omega)=0$, for all $(t,\omega)$. Then
$\lim_{n\to\infty}\bigg{\|}\int_{0}^{T}H^{n}_{s-}{\mathrm{d}}Z_{s}\bigg{\|}_{{\mathcal{H}}^{r}}=0.$
Next, we impose the compatibility condition for filtrations, in the following sense.
Assumption 2.6 (Compatibility).
The sub-filtration ${\mathbb{G}}\subset{\mathbb{F}}$ satisfies the compatibility condition, if for each $0\leq t\leq T$, ${\mathcal{F}}_{t}$ and ${\mathcal{G}}_{T}$ are conditionally independent given ${\mathcal{G}}_{t}$, expressed as ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}\big{|}{\mathcal{G}}_{t}.$
Remark 2.7.
There are a number of alternative characterizations for the compatibility condition, such as the immersion or the H-hypothesis, in the literature of filtration enlargement and shrinkage (see for instance [8, 26, 38]). Given two filtrations ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}$ and ${\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]}$ defined on the same space, with ${\mathbb{G}}\subset{\mathbb{F}}$, the following are known to be equivalent ([8, Theorem 3]):
  • ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$, for all $0\leq t\leq T.$ (Compatibility condition).
  • Every square integrable $\mathbb{G}$}-martingale is an $\mathbb{F}$}-martingale. (H-hypothesis).
  • ${\mathbb{E}}[Z|{\mathcal{G}}_{t}]$ = ${\mathbb{E}}[Z|{\mathcal{F}}_{t}]$ a.s., for each $t\in[0,T]$ and each bounded ${\mathcal{F}}_{T}$-measurable random variable $Z$.
In the context of mean-field games, this compatibility assumption ensures a weak closure of adapted processes, as first noted in [15, Lemma 3.11] and stated more explicitly in [6, Theorem 5.4].
Moreover, as shown in the following lemma, this compatibility condition allows us to define a càdlàg version of $\mu_{t}$ by taking $\mu_{t}=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})$, which will be used throughout the paper.
Lemma 2.8.
Under Assumption 2.6, $\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})$ almost surely.
Proof.
It suffices to show that for any bounded ${\mathcal{G}}_{T}$- measurable random variable $A$, ${\mathbb{E}}\big{[}\varphi(\bm{X}_{t})A\big{]}={\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(\bm{X}_{t})|{\mathcal{G}}_{t}\big{]}A\Big{]}.$ This is guaranteed by ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$ and
${\mathbb{E}}\big{[}\varphi(X_{t})A\big{]}$ $={\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(X_{t})A|{\mathcal{G}}_{t}\big{]}\Big{]}={\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(X_{t})|{\mathcal{G}}_{t}\big{]}{\mathbb{E}}\big{[}A|{\mathcal{G}}_{t}\big{]}\Big{]}$
$={\mathbb{E}}\Big{[}{\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(X_{t})|{\mathcal{G}}_{t}\big{]}A\big{|}{\mathcal{G}}_{t}\Big{]}\Big{]}={\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(X_{t})|{\mathcal{G}}_{t}\big{]}A\Big{]}.$
Finally, we will assume $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ according to the following definition.
Definition 2.9 ($C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$).
For $p\geq 2$, a function $\Phi(\mu,\bm{y})$ is $C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ if for any $\mu\in\mathcal{P}_{p}({\mathbb{R}}^{d})$ and $\bm{y}\in{\mathbb{R}}^{d}$, there exists a continuous mapping $(\mu,\bm{y},\bm{x}_{1})\to\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, as well as a continuous mapping $(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})\to\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})$ that is symmetric in its two arguments $(\bm{x}_{1},\bm{x}_{2})$ and with the following properties:
  • (Continuous differentiability) $\nabla_{\bm{y}}\Phi(\mu,\bm{y})$, $\nabla_{\bm{y}}^{2}\Phi(\mu,\bm{y})$, $\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$,
    $\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$ $\nabla^{2}_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\nabla_{\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})$ all exist and are continuous for all $\bm{y},\bm{x}_{1},\bm{x}_{2}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$.
  • (Uniform polynomial-growth) there exists a constant $c>0$ such that for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$,
    $\Big{|}\nabla_{\bm{y}}\Phi\Big{|}\leq c\big{(}1+|\bm{y}|^{p-1}\big{)},\quad\Big{|}\nabla^{2}_{\bm{y}}\Phi\Big{|}$ $\leq c\big{(}1+|\bm{y}|^{p-2}\big{)},$
    $\bigg{|}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\bigg{|}+\bigg{|}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-1}+|\bm{y}|^{p-1}\big{)},$
    $\bigg{|}\nabla^{2}_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\bigg{|}+\bigg{|}\nabla_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{y}|^{{p-2}}\big{)},$
    $\bigg{|}\nabla_{\bm{x}_{1},\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{x}_{2}|^{p-2}+|\bm{y}|^{p-2}\big{)}.$
  • (Fundamental theorem of calculus) for everything $\mu,\nu\in{\mathcal{P}}_{2}({\mathbb{R}}^{d})$,
    $\Phi(\mu,\bm{y})$- $\Phi(\nu,\bm{y})=\int_{0}^{1}\int_{{\mathbb{R}}^{d}}\frac{\delta\Phi}{\delta\mu}\big{(}\lambda\mu+(1-\lambda)\nu,\bm{y},\bm{x}_{1}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1}){\mathrm{d}}\lambda,$
    $\Phi(\mu,\bm{y})$- $\Phi(\nu,\bm{y})-\int_{0}^{1}\int_{{\mathbb{R}}^{d}}\frac{\delta\Phi}{\delta\mu}\big{(}\nu,\bm{y},\bm{x}_{1}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1})$
    $=\int_{0}^{1}\int_{0}^{t}\int_{{\mathbb{R}}^{d}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}s\mu+(1-s)\nu,\bm{y},\bm{x}_{1},\bm{x}_{2}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1})(\mu-\nu)({\mathrm{d}}\bm{x}_{2}){\mathrm{d}}s{\mathrm{d}}t.$
Here $\frac{\delta\Phi}{\delta\mu}$ and $\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}$ are respectively (a version of) the first order and the second order linear derivatives of $\Phi$.
Remark 2.10.
Note that the linear derivative is in the sense of [21], which is properly defined up to an additive constant. This form of linear derivative is shown to be appropriate for analyzing the flow of measures on semimartingales [21, 31].
2.2. Main result
Now we are ready to state the main result.
Theorem 2.11 (Itô’s formula for flows of conditional measure on semimartingales).
[…]
$\Phi(\mu_{t},\bm{Y}_{t})-\Phi(\mu_{0},\bm{Y}_{0})$ (2.3)
$=$ $\;\overline{{\mathbb{E}}}\bigg{[}\int_{0+}^{t}\bigg{(}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}$
$\qquad\quad+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{s}+\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]^{c}_{s}\bigg{)}$
$\quad\quad+\sum_{0<s\leq t}\bigg{(}\frac{1}{2}\bigg{(}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$
$\qquad\qquad\qquad-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}\bigg{)}$
$\qquad\qquad\qquad+\bigg{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{)}\cdot\Delta\bm{Y}_{s}$
$\quad\qquad\qquad+\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{)}\mathbbm{1}_{\{\mu_{s}=\mu_{s-}\}}\bigg{|}{\mathcal{F}}\bigg{]}$
$+\int_{0+}^{t}\nabla_{\bm{y}}\Phi(\mu_{s-},\bm{Y}_{s-})\cdot{\mathrm{d}}(\bm{Y})^{c}_{s}+\frac{1}{2}\int_{0+}^{t}\nabla_{\bm{y}\bm{y}}^{2}\Phi(\mu_{s-},\bm{Y}_{s-}):{\mathrm{d}}[\bm{Y},\bm{Y}]^{c}_{s}$
$+\sum_{0<s\leq t}\Big{(}\Phi(\mu_{s},\bm{Y}_{s})-\Phi(\mu_{s-},\bm{Y}_{s-})\Big{)},$
for all $t\in[0,T]$, where $\bm{X}^{\prime},\bm{X}^{\prime\prime}$ are two conditionally independent copies of $\bm{X}$ given the sub $\sigma$-algebra ${\mathcal{G}}_{T}$ defined in some enlarged probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$, and $\overline{{\mathbb{E}}}[\cdot|{\mathcal{F}}]$ is the conditional expectation given $\mathcal{F}$} in the enlarged probability space.
Remark 2.12.
To ensure that the enlarged space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}},\overline{{\mathbb{P}}})$ is properly defined such that $\mu_{t},\bm{X}_{t},\bm{Y}_{t}$ and $\mathcal{F}$} are naturally extended to this enlarged space and $\bm{X}^{\prime}$$\bm{X}^{\prime\prime}$ are well-defined, one needs the precise notion of conditionally independent copies of stochastic processes. In addition, one needs to show the existence of these conditionally independent copies in Theorem 2.11. These will be studied in the next section and specified in Theorem 3.1 and Corollary 3.2.
Remark 2.13.
Theorem 2.11 includes the following Itô’s formulas as special cases:
  • The Itô’s formula for mean-field games and control problems with common noise developed in [9, 14]: take $X_{t}$ as an Itô’s process driven by both idiosyncratic noise $W$ and common noise $W^{0}$, let the common filtration be the filtration generated by $W_{0}$.
  • The Itô’s formula for flows of measures associated with possibly discontinuous processes without common noise studied in [31, 53]: set $X_{t}$ as a general semimartingale and choose $\mathcal{G}$} as the trivial $\sigma$-algebra.
  • The Itô’s formula for partially observed diffusions as in [4]: let $X_{t}$ be an Itô’s process and letting $\mathcal{G}$} be the filtration generated by the observation process. For more details, see Section 6 for an extension of this problem to jump-diffusion processes.
3. Conditionally Independent Copy of Stochastic Process
3.1. Why conditionally independent copy?
One of the key components in Theorem 2.11 is the notion of conditionally independent copies of stochastic processes. We now illustrate this concept through several examples in the context of flow of conditional laws for semimartingales and its role in the derivation of Itô’s formula.
Let us first consider a real-valued stochastic process $X_{t}$ of the form
${\mathrm{d}}X_{t}=b_{t}{\mathrm{d}}t+\sigma_{t}{\mathrm{d}}W_{t}+\sigma^{0}_{t}{\mathrm{d}}W^{0}_{t},$
where $W,W^{0}$ are two independent real valued Brownian motions, ${\mathbb{G}}=\{{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is the filtration generated by $W^{0}$, ${\mathcal{G}}={\mathcal{G}}_{T}=\sigma(W^{0}_{u},u\leq T)$, and $\sigma_{t}$ and $\sigma_{t}^{0}$ are appropriate functions to ensure the well-definedness of the above SDE. By the classical Itô’s formula, the dynamic of $g(X_{t})$ for a smooth function $g$ is given by
${\mathrm{d}}g(X_{t})=b_{t}g^{\prime}(X_{t})+\frac{1}{2}(\sigma_{t}^{2}+(\sigma^{0}_{t})^{2})g^{\prime\prime}(X_{t}){\mathrm{d}}t+\sigma_{t}g^{\prime}(X_{t}){\mathrm{d}}W_{t}+\sigma^{0}_{t}g^{\prime}(X_{t}){\mathrm{d}}W^{0}_{t},$
and consequently the dynamic of $Z_{t}={\mathbb{E}}[g(X_{t})|{\mathcal{G}}]$ is given by
${\mathrm{d}}Z_{t}={\mathbb{E}}\bigg{[}b_{t}g^{\prime}(X_{t})+\frac{1}{2}(\sigma_{t}^{2}+(\sigma^{0}_{t})^{2})g^{\prime\prime}(X_{t})\Big{|}{\mathcal{G}}\bigg{]}{\mathrm{d}}t+{\mathbb{E}}\big{[}\sigma^{0}_{t}g^{\prime}(X_{t})|{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{t}.$ (3.1)
The above equation follows essentially from the fact that for progressively measurable process $(\eta_{t})_{t\in[0,T]},$
${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}W_{s}\Big{|}{\mathcal{G}}\bigg{]}=0,\qquad{\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}W^{0}_{s}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}W^{0}_{s},\text{ for all }t\in[0,T].$ (3.2)
(See [47, Lemma B.1] for more details).
From (3.1) one can calculate the quadratic variation of $Z_{t}$:
${\mathrm{d}}\langle Z\rangle_{t}={\mathbb{E}}\big{[}\sigma^{0}_{t}g^{\prime}(X_{t})|{\mathcal{G}}\big{]}^{2}{\mathrm{d}}t={\mathbb{E}}[\sigma^{0}_{t}g^{\prime}(X_{t})\widetilde{\sigma}^{0}_{t}g^{\prime}(\widetilde{X}_{t})|{\mathcal{G}}]{\mathrm{d}}t,$
where $\widetilde{\sigma}_{t}^{0},\widetilde{X}_{t}$ are respectively conditionally independent copies of $\sigma^{0}_{t}$ and $X_{t}$ given $\mathcal{G}$}, as formally introduced in [14]. Moreover, one can derive for a smooth function $f$ the Itô’s formula for $f(Z_{t})$, which depends on the precise expression of
$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s}=$ $\int_{0}^{t}f^{\prime}(Z_{s})\bigg{(}{\mathbb{E}}\Big{[}b_{s}g^{\prime}(X_{s})+\frac{1}{2}(\sigma_{s}^{2}+(\sigma^{0}_{s})^{2})g^{\prime\prime}(X_{s})\Big{|}{\mathcal{G}}\Big{]}{\mathrm{d}}s+{\mathbb{E}}\big{[}\sigma^{0}_{t}g^{\prime}(X_{s})|{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{s}\bigg{)}$
$=\int_{0}^{t}\bigg{(}{\mathbb{E}}$ $\Big{[}f^{\prime}(Z_{s})\Big{(}b_{s}g^{\prime}(X_{s})+\frac{1}{2}(\sigma_{s}^{2}+(\sigma^{0}_{s})^{2})g^{\prime\prime}(X_{s})\Big{)}\Big{|}{\mathcal{G}}\Big{]}{\mathrm{d}}s+{\mathbb{E}}\big{[}f^{\prime}(Z_{s})\sigma^{0}_{t}g^{\prime}(X_{s})|{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{s}\bigg{)}$
and
$\int_{0}^{t}f^{\prime\prime}(Z_{s}){\mathrm{d}}\langle Z\rangle_{s}$ $=\int_{0}^{t}f^{\prime\prime}(Z_{s})\cdot\frac{1}{2}{\mathbb{E}}\big{[}\sigma^{0}_{s}g^{\prime}(X_{s})\widetilde{\sigma}^{0}_{s}g^{\prime}(\widetilde{X}_{s})\big{|}{\mathcal{G}}_{T}\big{]}{\mathrm{d}}s$
$=\frac{1}{2}\int_{0}^{t}{\mathbb{E}}\big{[}f^{\prime\prime}(Z_{s})\sigma^{0}_{s}g^{\prime}(X_{s})\widetilde{\sigma}^{0}_{s}g^{\prime}(\widetilde{X}_{s})\big{|}{\mathcal{G}}_{T}\big{]}{\mathrm{d}}s.$
However, there is an issue when one considers a general continuous semimartingale $X$ and a general sub $\sigma$-algebra $\mathcal{G}$}, where the Itô’s formula takes the form:
${\mathrm{d}}g(X_{t})=g^{\prime}(X_{t}){\mathrm{d}}X_{t}+\frac{1}{2}g^{\prime\prime}(X_{t}){\mathrm{d}}\langle X\rangle_{t}.$ (3.3)
Indeed, in order to characterize the dynamics of $Z_{t}={\mathbb{E}}[g(X_{t})|{\mathcal{G}}]$, one needs an appropriate expression for ${\mathrm{d}}Z_{t}$, ${\mathrm{d}}\langle Z\rangle_{t}$, and more importantly for
[…]
One naive guess of the extension of (3.2) to this general case is
${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}{\mathbb{E}}[X_{s}|{\mathcal{G}}].$ (3.5)
Unfortunately this guess does not hold in general. For instance, take $X_{t}$ from (3.1) with $b_{t}=0$. By (3.2), the left hand side of (3.5) becomes
${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}\Big{|}{\mathcal{G}}\bigg{]}={\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}\big{(}\sigma_{s}{\mathrm{d}}W_{s}+\sigma^{0}_{s}{\mathrm{d}}W^{0}_{s}\big{)}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}\big{[}\eta_{s}\sigma^{0}_{s}\big{|}{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{s},$
while the right hand side is
$\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}X^{\mathcal{G}}_{s}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathbb{E}}[\sigma^{0}_{s}|{\mathcal{G}}]{\mathrm{d}}W^{0}_{s},$
and (3.5) holds if and only if $\eta_{s}$ and $\sigma_{s}^{0}$ are conditionally uncorrelated, given $\mathcal{G}$}.
[…]
$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s}$ $=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}{\mathbb{E}}\big{[}g(X_{s})|{\mathcal{G}}\big{]}=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}\widetilde{\mathbb{E}}\big{[}g(X^{(1)}_{s})\big{]}=\widetilde{\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}g(X^{(1)}_{s})\bigg{]}$
$=\widetilde{\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s})\Big{(}g^{\prime}(X^{(1)}_{s}){\mathrm{d}}X^{(1)}_{s}+\frac{1}{2}g^{\prime\prime}(X^{(1)}_{s}){\mathrm{d}}\langle X^{(1)}\rangle_{s}\Big{)}\bigg{]},$
and
$\int_{0}^{t}f^{\prime\prime}(Z_{s}){\mathrm{d}}\langle Z\rangle_{s}$ $=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}\langle{\mathbb{E}}\big{[}g(X_{\cdot})|{\mathcal{G}}\big{]}\rangle_{s}=\widetilde{\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}\langle g(X^{(1)}_{\cdot}),g(X^{(2)}_{\cdot})\rangle_{s}\bigg{]}$
$=\widetilde{\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime\prime}(Z_{s})g^{\prime}(X^{(1)}_{s})g^{\prime}(X^{(2)}_{s}){\mathrm{d}}\langle X^{(1)},X^{(2)}\rangle_{s}\bigg{]},$
where $\widetilde{\mathbb{E}}$ denotes the expectation on $X^{(1)}$ and $X^{(2)}$ while keeping the information that is $\mathcal{G}$}-measurable. With these expressions, one can analyze (3.4) for a general semimartingale $X$, as will be detailed in (1) and (2) of Theorem 3.5.
Next, we will provide the construction of conditionally independent copies for the general case of a semimartingale $\bm{X}$ and an arbitrary sub $\sigma$-algebra $\mathcal{G}$} which is not necessarily associated with any random variable. As emphasized earlier, this construction is necessary for establishing Itô’s formula with the flow of conditional measures on semimartigales.
3.2. Construction of conditionally independent copies of stochastic process
Given a probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$, consider the extended probability space $\overline{\Omega}=\Omega^{n+1}$ defined by
$\overline{\Omega}=\Omega^{n+1}=\{(\omega_{0},\omega_{1},\cdots,\omega_{n})|\omega_{0},\omega_{1},\cdots,\omega_{n}\in\Omega\},$ (3.6)
and the extended $\sigma$-algebra $\overline{{\mathcal{F}}}$ by
$\overline{{\mathcal{F}}}=\sigma\{A_{0}\times A_{1}\times\cdots\times A_{n}:A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}\}.$ (3.7)
To ensure that these copies are conditionally independent, let us define the probability measure $\overline{{\mathbb{P}}}$ as follows: for $A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}$,
$\overline{{\mathbb{P}}}\big{(}A_{0}\times A_{1}\times\cdots\times A_{n}\big{)}:={\mathbb{E}}\bigg{[}\mathbbm{1}_{A_{0}}\prod_{i=1}^{n}{\mathbb{P}}(A_{i}|{\mathcal{G}})\bigg{]}.$ (3.8)
This measure is well defined since the set $\{A_{0}\times A_{1}\times\cdots\times A_{n}\}$ is a $\pi$-system. Finally, we define the completion of the above extended probability space and still use $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ to denote the completed probability space. Further, define $\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$ for the extension of $\mathcal{F}$} and $\mathcal{G}$} into this enlarged probability space as
[…]
Then we have:
Theorem 3.1.
Given a probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$ and a sub $\sigma$-algebra $\mathcal{G}$$\subset{\mathcal{F}}$, there exists an enlarged completed probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ satisfying (3.6), (3.7) and (3.8); and for any random variable $R$ taking values in a standard Borel space $(E,\mathcal{E})$, there exist an $\widetilde{R}$ and $n$ conditionally independent copies $\{R^{(i)}\}_{i=1,2,..n}$ given $\widetilde{\mathcal{G}}$ such that
$\widetilde{R}(\omega_{0},\omega_{1},\cdots,\omega_{n})=R(\omega_{0}),\;R^{(i)}(\omega_{0},\omega_{1},\cdots,\omega_{n})=R(\omega_{i}),$ (3.10)
with $\widetilde{R}$ and $\{R^{(i)}\}_{i=1,2,..n}$ satisfying the following:
  • Given $\widetilde{\mathcal{G}}$, $\widetilde{R}$ and $\big{\{}R^{(i)}\big{\}}_{i=1,2,..n}$ are conditionally independent.
  • Given $\widetilde{\mathcal{F}}$, $\widetilde{R}$ and $\big{\{}R^{(i)}\big{\}}_{i=1,2,..n}$ are conditionally independent.
  • $\big{(}R\big{|}{\mathcal{G}}\big{)}(\omega_{0})=Law$$\big{(}\widetilde{R}\big{|}\widetilde{\mathcal{G}}\big{)}(\overline{\omega})=$$\big{(}R^{(i)}\big{|}{\widetilde{\mathcal{G}}}\big{)}(\overline{\omega})=Law$$\big{(}R^{(i)}\big{|}{\widetilde{\mathcal{F}}}\big{)}(\overline{\omega})$, for all $i=1,2,..,n$, a.s.,
for $\overline{\omega}=(\omega_{0},,\omega_{1},\cdots,\omega_{n})\in\overline{\Omega}$.
Note that when $\mathcal{G}$} is a trivial $\sigma$-algebra $\{\emptyset,\Omega\}$, then the above theorem reduces to the standard construction of independent copies of stochastic processes.
3.3. Construction of independent copies of semimartingale
In the context of semimartingales, we can similarly define a corresponding version of the conditionally independent copy as follows. Take the completed filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$ with a sub-filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$ satisfying ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$, define a complete extended probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ satisfying (3.6), (3.7) and (3.8) by setting the sub $\sigma$-algebra ${\mathcal{G}}={\mathcal{G}}_{T}$. Define $\overline{{\mathbb{F}}}:=(\overline{{\mathcal{F}}}_{t})_{t\in(0,T)}$ by
$\overline{{\mathcal{F}}}_{t}:=\text{completion of }\sigma\{A_{0}\times A_{1}\times\cdots\times A_{n}:A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}_{t}\},$ (3.11)
and define $\widetilde{\mathbb{F}}:=(\widetilde{\mathcal{F}}_{t})_{t\in(0,T)}$ and $\widetilde{\mathbb{G}}:=(\widetilde{\mathcal{G}}_{t})_{t\in(0,T)}$ respectively for the extension of $\mathbb{F}$} and $\mathbb{G}$} in this enlarged probability space as
[…]
With this construction, we have
Corollary 3.2.
[…]
$(\bm{X}^{\prime},\bm{Y}^{\prime})(\omega_{0},\omega_{1},\cdots,\omega_{n})=(\bm{X},\bm{Y})(\omega_{1}),\;(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})(\omega_{0},\omega_{1},\cdots,\omega_{n})=(\bm{X},\bm{Y})(\omega_{2}),$
$(\widetilde{\bm{X}},\widetilde{\bm{Y}})(\omega_{0},\omega_{1},\cdots,\omega_{n})=(\bm{X},\bm{Y})(\omega_{0}),$
[…]
$\begin{cases}\text{Law}\big{(}\widetilde{\bm{X}},\widetilde{\bm{Y}}\big{|}\widetilde{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime},\bm{Y}^{\prime}\big{|}\widetilde{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime},\bm{Y}^{\prime}\big{|}\widetilde{\mathcal{F}}\big{)}\\ \qquad=\text{Law}\big{(}\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime}\big{|}\widetilde{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime}\big{|}\widetilde{\mathcal{F}}\big{)}\text{ a.s.,}\\ (\widetilde{\bm{X}},\widetilde{\bm{Y}}),(\bm{X}^{\prime},\bm{Y}^{\prime}),(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})\text{ are conditionally independent, given }{\mathcal{G}}\ or\ {\mathcal{F}}.\par\par\end{cases}$ (3.13)
Furthermore, if the sub $\sigma$-algebra ${\mathcal{G}}={\mathcal{G}}_{T}$, and ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}\big{|}{\mathcal{G}}_{t},$ for all $0\leq t\leq T$, then the sub-filtration $\widetilde{\mathcal{G}}\subset\widetilde{\mathcal{F}}$ satisfies $\widetilde{\mathcal{F}}_{t}\perp\!\!\!\perp\widetilde{\mathcal{G}}_{T}\big{|}\widetilde{\mathcal{G}}_{t},$ for all $0\leq t\leq T$.
[…]
$\widetilde{\mu}_{t}(\omega_{0},\omega_{1},\cdots,\omega_{n})=\text{Law}(\widetilde{\bm{X}_{t}}\big{|}\widetilde{\mathcal{G}}_{T})(\omega_{0},\omega_{1},\cdots,\omega_{n})=\text{Law}(\bm{X}_{t}\big{|}\widetilde{\mathcal{G}}_{T})(\omega_{0})=\mu_{t}(\omega_{0}),$
which in turn allows for using $\mu_{t}$ for $\widetilde{\mu}_{t}$ without any ambiguity. Therefore, from now on we sometimes will simply use the original notation to denote the natural extensions.
3.4. Examples
To get some intuition for this notion of conditionally independent copies of stochastic processes, let us see some special examples of $\mathcal{F}$} and the sub $\sigma$-algebra $\mathcal{G}$}.
Example 1. Let $A,B$ be two independent random variables with ${\mathcal{F}}=\sigma(A,B)$ and ${\mathcal{G}}=\sigma(B)$. One can enlarge the probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ to include $\{(A^{(i)},B^{(i)})\}_{i=1,2,..,n}$ satisfying (3.6), (3.7),(3.8) and (3.10). Moreover,
  1. (1)
    $\{(A^{(i)},B^{(i)})\}_{i=1,2,..,n}$ have the same distribution as $(A,B)$.
  2. (2)
    $B^{(i)}=B$ almost surely.
  3. (3)
    $\overline{{\mathcal{F}}}=\sigma\big{(}A,B,\{A^{(i)},B^{(i)}\}_{i=1,2,..,n}\big{)}=\sigma\big{(}A,B,\{A^{(i)}\}_{i=1,2,..,n}\big{)}$.
  4. (4)
    $\{A^{(i)}\}_{i=1,2,..,n}$ and $A,B$ are independent.
Consequently, for any $X=f(A,B)$, the conditionally independent copies of $X$ are $X^{(i)}=f(A^{(i)},B^{(i)})=f(A^{(i)},B),\ \ i=1,\cdots,n.$
Example 2. Let $X$ satisfy the following stochastic differential equation
${\mathrm{d}}X_{t}=b(X_{t}){\mathrm{d}}t+\sigma(X_{t}){\mathrm{d}}W_{t}+\sigma^{(0)}(X_{t}){\mathrm{d}}W^{(0)}_{t}$,
where $W,W^{(0)}$ are two independent Brownian motions, and $b$, $\sigma$, and $\sigma^{0}$ satisfy appropriate continuity conditions. Let ${\mathcal{F}}_{t}=\sigma\big{(}\{W_{s},W^{(0)}_{s}\}_{s\leq t}\big{)}$ and the sub-filtration ${\mathcal{G}}_{t}=\sigma\big{(}\{W^{(0)}_{s}\}_{s\leq t}\big{)}$, then one can enlarge the probability space to include independent Brownian motions $\big{\{}(X^{\prime},W^{\prime},{W^{(0)}}^{\prime})\big{\}}_{i=1,2,..,N}$ satisfying (3.6), (3.7),(3.8), (3.10) and (3.11). Moreover,
  1. (1)
    $X^{\prime}$ satisfies the following stochastic differential equation
    ${\mathrm{d}}X^{\prime}_{t}=b(X^{\prime}_{t}){\mathrm{d}}t+\sigma(X^{\prime}_{t}){\mathrm{d}}W^{\prime}_{t}+\sigma^{(0)}(X^{\prime}_{t}){\mathrm{d}}(W^{(0)}_{t})^{\prime}$.
  2. (2)
    $W^{(0)}=(W^{(0)})^{\prime}$ almost surely.
  3. (3)
    $\overline{{\mathcal{F}}}_{t}=\sigma\big{(}\{W_{s},W^{(0)}_{s},W^{\prime}_{s},(W^{(0}_{s})^{\prime}\}_{s\leq t}\big{)}=\sigma\big{(}\{W_{s},W^{(0}_{s},W^{\prime}_{s}\}_{s\leq t}\big{)}$
  4. (4)
    $W,W^{\prime},W^{(0)}$ are independent Brownian motions.
Consequently, $X^{\prime}$ is one conditionally independent copy of $X$ given $\mathcal{G}$}.
There are more examples of conditionally independent copies of stochastic processes in the literature of mean-field games. For instance, when defining a semimartingale by stochastic differential equations involving both idiosyncratic and common noise, conditionally independent copies are formally formulated in [14] as $\overline{\Omega}=\Omega_{0}\times\Omega_{1}\times\widetilde{\Omega}_{1}$, where the original probability space follows the structure $\Omega=\Omega_{0}\times\Omega_{1}$, and $\widetilde{\Omega}_{1}$ represents the independent copy of the probability space $\Omega_{1}$ ([14, Theorem 4.14, Volume II]).
3.5. Properties of the conditionally independent copies of semimartingales
Next, we study properties of these conditionally independent copies of semimartingales, and explore the relation between the conditional expectation of their stochastic integrals and stochastic integrals with respect to the conditional law of semimartingales. For notational simplicity, we will present the results for the one-dimensional case and their corresponding multi-dimensional cases can be obtained with appropriate adaptations.
First, by the compatibility condition and the classical results regarding filtration enlargement, we have:
Proposition 3.3.
Given an $\mathbb{R}$}-valued $\mathbb{F}$}-semimartingales $\{X_{t}\}_{t\in[0,T]}$ on a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$. Let $X^{\prime}$ and $X^{\prime\prime}$ be its conditionally independent copies defined in Proposition 3.2 in the enlarged filtered probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}}=(\overline{{\mathcal{F}}}_{t})_{t\in[0,T]},\overline{{\mathbb{P}}})$ satisfying (3.6), (3.8) and (3.11). Then,
  1. (1)
    $X$, $^{\prime}, X′′X^{\prime\prime}$ are $\overline{{\mathbb{F}}}$-semimartingales.
  2. (2)
    Suppose ${\mathbb{E}}[|X_{t}|]<\infty$ for all $t\in[0,T]$, and define
    $X^{\mathcal{G}}_{t}:=\overline{{\mathbb{E}}}\big{[}X_{t}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{{\mathbb{E}}}\big{[}X_{t}\big{|}{\mathcal{G}}_{t}\big{]}.$
    1. (a)
      If $\{X_{t}\}_{t\in[0,T]}$ is a finite variation process and ${\mathbb{E}}\big{[}\big{(}\int_{0}^{T}|{\mathrm{d}}X_{s}|\big{)}^{p}\big{]}<\infty$ for $p\geq 1$, then $\{X^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is also a finite variation process, and
      $\overline{{\mathbb{E}}}\bigg{[}\bigg{(}\int_{0}^{T}\big{|}{\mathrm{d}}X^{\mathcal{G}}_{s}\big{|}\bigg{)}^{p}\bigg{]}<\infty.$
    2. (b)
      $\{X^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is a semimartingale with respect to $\overline{{\mathbb{F}}}$, $\mathbb{F}$}, and $\mathbb{G}$}.
Remark 3.4.
Proposition 3.3 shows that semimartingales remain semimartingales under the enlarged filtration $\overline{{\mathbb{F}}}$. This is closely related to the so called $H^{\prime}$-hypothesis where any $\mathbb{F}$}-martingale is an $\overline{{\mathbb{F}}}$-semimartingale with the enlargement filtration from $\mathbb{F}$} to $\overline{{\mathbb{F}}}$ (see for instance [3, 29]). Here we have obtained a slightly stronger result that any $\mathbb{F}$}-semimartingale $X$ and its conditionally independent copy remain $\overline{{\mathbb{F}}}$-semimartingales; meanwhile the $\mathbb{G}$}-semimartingale $X^{\mathcal{G}}$ (the projection taken with respect to the $\sigma$-algebra $\mathcal{G}$} as in [8]) remains $\overline{{\mathbb{F}}}$-semimartingale. Despite the apparent similarity, we are unable to locate exactly the same statement, and will provide the proof for completeness.
Theorem 3.5.
Give two $\mathbb{R}$}-valued $\mathbb{F}$}-semimartingales $\{X_{t}\}_{t\in[0,T]}$ and $\{Y_{t}\}_{t\in[0,T]}$ on a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$. Let $(X^{\prime},Y^{\prime})$ and $(X^{\prime\prime},Y^{\prime\prime})$ be their conditionally independent copies defined in Proposition 3.2 in the enlarged filtered probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}}=({\mathcal{F}}_{t})_{t\in[0,T]},\overline{{\mathbb{P}}})$ satisfying (3.6), (3.8) and (3.11). Suppose ${\mathbb{E}}[|X_{t}|]<\infty$ and ${\mathbb{E}}[|Y_{t}|]<\infty$ for all $t\in[0,T]$.
  1. (1)
    If $\{Z_{t}\}_{t\in[0,T]}$ is an $\mathbb{R}$}-valued $\mathbb{F}$}-adapted càdlàg process in the filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$, and $\|X\|_{\mathcal{H}_{p}}+\|Z\|_{\mathcal{H}_{q}}<\infty$ with $\frac{1}{p}+\frac{1}{q}=1$, then
    $\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{|}{\mathcal{F}}\bigg{]}.$
  2. (2)
    If $\{Z_{t}\}_{t\in[0,T]}$ is an $\mathbb{R}$}-valued $\mathbb{F}$}-adapted càdlàg process in a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$ and $\|X\|_{\mathcal{H}_{p}}+\|Y\|_{\mathcal{H}_{q}}+\|Z\|_{\mathcal{H}_{r}}<\infty$ with $\frac{1}{p}+\frac{1}{q}+\frac{1}{r}=1$, then
    $\int_{0}^{t}Z_{s-}{\mathrm{d}}\big{[}X^{\mathcal{G}},Y^{\mathcal{G}}\big{]}_{s}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}\big{[}X^{\prime\prime},Y^{\prime}\big{]}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
    $\int_{0}^{t}Z_{s}{\mathrm{d}}\big{[}X^{\mathcal{G}},Y\big{]}_{s}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z_{s}{\mathrm{d}}\big{[}X^{\prime},Y\big{]}_{s}\bigg{|}{\mathcal{F}}\bigg{]}.$
Remark 3.6.
Statements (1) and (2) in Theorem 3.5 are the key equations for deriving the Itô’s formula (Theorem 2.11. They ensure crucially a) the equivalence between the stochastic integral of the conditional expectation of a semimartingale and the conditional expectation of the Itô’s integral of its conditionally independent copy, and b) the equivalence between the quadratic variation of the conditional expectation of a semimartingale and the conditional expectation of the quadratic variation of its two conditionally independent copies.
The proof of Theorem 3.5 is presented in Section 5.3.
4. Proof of Main Result (Theorem 2.11)
Having established necessary properties of conditionally independent copies of semimartingales, we are now ready to prove Theorem 2.11, the Itô’s formula for flow of conditional laws on semimartingales, through several steps. First, we will derive the formula for cylindrical functions; we will then show that cylindrical functions are dense in $C^{2,2}({\mathbb{R}}^{d}\times\mathcal{P}_{p}({\mathbb{R}}^{d}))$; finally, with appropriate localization technique and dominated convergence theorem, we will finish the proof.
4.1. Step 1: Ito’s formula for cylindrical functions.
First, let us recall the cylindrical functions (see for instance ([23]).
Definition 4.1 ($C^{2,2}$ cylindrical functions).
A function $\Phi:{\mathcal{P}}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ is a $C^{2,2}$ cylindrical function if
$\Phi(\mu,\bm{y})=f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)},$
where $f\in C^{2}({\mathbb{R}}^{n+d})$ and $g=(g^{(1)},..,g^{(n)})$, with $g^{(i)}\in C^{2}_{b}({\mathbb{R}}^{d})$.
We will first show that for smooth cylindrical functions $\Phi$, Theorem 2.11 holds. Indeed, for semimartingales $\{\bm{X}_{t}\}_{t\in[0,T]}$ and $g^{(i)}\in C_{b}^{2}({\mathbb{R}}^{d})$ with $i=1,2,..,n$, by the classical Itô’s formula [51, Theorem 33, Chapter II],
$g^{(i)}(\bm{X}_{t})-g^{(i)}(\bm{X}_{0})=$ $\int_{0+}^{t}\nabla g^{(i)}(\bm{X}_{s-})\cdot{\mathrm{d}}\bm{X}^{c}_{s}+\frac{1}{2}\int_{0+}^{t}\nabla^{2}g^{(i)}(\bm{X}_{s-}):{\mathrm{d}}[\bm{X},\bm{X}]^{c}_{s}$
$\;\;+\sum_{0<s\leq t}\Big{\{}g^{(i)}(\bm{X}_{s})-g^{(i)}(\bm{X}_{s-})\Big{\}},\text{ for }t\in[0,T],$
and $\{g^{(i)}(\bm{X}_{t})\}_{t\in[0,T]}$ for $i=1,..,n$ are semimartingales. Taking conditional expectation given ${\mathcal{G}}_{T}$, and recall that $\mu_{t}=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})$, we have
$\big{\langle}\mu_{t},g^{(i)}\big{\rangle}-\big{\langle}\mu_{0},g^{(i)}\big{\rangle}=$ $\;{\mathbb{E}}\bigg{[}\int_{0+}^{t}\nabla g^{(i)}(\bm{X}_{s-})\cdot{\mathrm{d}}\bm{X}^{c}_{s}+\frac{1}{2}\int_{0+}^{t}\nabla^{2}g^{(i)}(\bm{X}_{s-}):{\mathrm{d}}[\bm{X},\bm{X}]^{c}_{s}$ (4.1)
$\qquad+\sum_{0<s\leq t}\Big{\{}g^{(i)}(\bm{X}_{s})-g^{(i)}(\bm{X}_{s-})\Big{\}}\bigg{|}{\mathcal{G}}_{T}\bigg{]}.$
Define $\bm{Z}$ by
$\bm{Z}:=\Big{\{}\bm{Z}_{t}=\Big{(}\big{\langle}\mu_{t},g^{(1)}\big{\rangle},\big{\langle}\mu_{t},g^{(2)}\big{\rangle},..,\big{\langle}\mu_{t},g^{(n)}\big{\rangle}\Big{)}\Big{\}}_{t\in[0,T]},$
then by (2b) of Theorem 3.5, $\bm{Z}$ is an $\overline{{\mathbb{F}}}$-semimartingale.
Let us for the moment assume that $\bm{Y}_{t}$ is bounded. Then applying Itô’s formula to $f\big{(}\bm{Z},\bm{Y}\big{)}$ yields
$f(\bm{Z}_{t},\bm{Y}_{t})-f(\bm{Z}_{0},\bm{Y}_{0})$ (4.2)
$=$ $\underbrace{\int_{0+}^{t}\nabla_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})\cdot{\mathrm{d}}\bm{Z}_{s}}_{J_{1}}+\frac{1}{2}\underbrace{\int_{0+}^{t}\nabla_{\bm{z}\bm{z}}^{2}f(\bm{Z}_{s-},\bm{Y}_{s-}):{\mathrm{d}}[\bm{Z},\bm{Z}]_{s}+\int_{0+}^{t}\nabla_{\bm{z}\bm{y}}^{2}f(\bm{Z}_{s-},\bm{Y}_{s-}):{\mathrm{d}}[\bm{Z},\bm{Y}]_{s}}_{J_{2}}$
- $\underbrace{\sum_{0<s\leq t}\Big{\{}\nabla_{\bm{z}}f(\bm{Y}_{s-},\bm{Z}_{s-})\cdot\Delta\bm{Z}_{s}+\frac{1}{2}\nabla^{2}_{\bm{z}\bm{z}}f(\bm{Y}_{s-},\bm{Z}_{s-}):\big{(}\Delta\bm{Z}_{s}\Delta\bm{Z}_{s}^{\top}\big{)}+\nabla^{2}_{\bm{z}\bm{y}}f(\bm{Y}_{s-},\bm{Z}_{s-}):\big{(}\Delta\bm{Z}_{s}\Delta\bm{Y}_{s}^{\top}\big{)}\Big{\}}}_{J_{3}}$
$\quad+\int_{0+}^{t}\nabla_{\bm{y}}f(\bm{Z}_{s-},\bm{Y}_{s-})\cdot{\mathrm{d}}\bm{Y}^{c}_{s}+\frac{1}{2}\int_{0+}^{t}\nabla_{\bm{y}\bm{y}}^{2}f(\bm{Z}_{s-},\bm{Y}_{s-}):{\mathrm{d}}[\bm{Y},\bm{Y}]^{c}_{s}$
$\quad+\sum_{0<s\leq t}\Big{\{}f(\bm{Y}_{s},\bm{Z}_{s})-f(\bm{Y}_{s-},\bm{Z}_{s-})\Big{\}}.$
We will compute $J_{1},J_{2},$ and $J_{3}$ separately. First, note that both $g^{(i)}$ and $\bm{Y}_{t}$ are bounded, and $f\in C^{2}({\mathbb{R}}^{n+d})$, therefore $\nabla_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})$, $\nabla^{2}_{\bm{z}\bm{y}}f(\bm{Z}_{s-},\bm{Y}_{s-})$, and $\nabla^{2}_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})$ are bounded, and the conditions for applying (1) and (2) of Theorem 3.5 are satisfied.
For the conditionally independent copy $\bm{X}^{\prime}$, we have
$g^{(i)}(\bm{X}^{\prime}_{t})-g^{(i)}(\bm{X}^{\prime}_{0})=$ $\int_{0+}^{t}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}+\frac{1}{2}\int_{0+}^{t}\nabla^{2}g^{(i)}(\bm{X}_{s-}):{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}$
$\;\;+\sum_{0<s\leq t}\Big{\{}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\Big{\}},\text{ for }t\in[0,T].$
Therefore, by (1) of Theorem 3.5, ${\mathrm{d}}\bm{Z}_{s}={\mathrm{d}}\big{(}{\mathbb{E}}[g^{(1)}(\bm{X}_{s})|{\mathcal{G}}],\cdots,{\mathbb{E}}[g^{(n)}(\bm{X}_{s})|{\mathcal{G}}]\big{)}$, and
$J_{1}=\overline{{\mathbb{E}}}\Bigg{[}\int_{0+}^{t}$ $\sum_{i=1}^{n}\partial_{z_{i}}f(\bm{Z}_{s-},\bm{Y}_{s-})\bigg{(}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}+\frac{1}{2}\nabla^{2}g^{(i)}(\bm{X}^{\prime}_{s-}):{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}\bigg{)}$ (4.3)
$+\sum_{0<s\leq t}\sum_{i=1}^{n}\partial_{z_{i}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{\{}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\Big{\}}\bigg{|}{\mathcal{F}}\Bigg{]}.$
For $J_{2}$, we see that
$\big{[}$ $g^{(i)}(\bm{X}^{\prime}_{t}),g^{(j)}(\bm{X}^{\prime\prime}_{t})\big{]}_{t}=\big{[}(g^{(i)}(\bm{X}^{\prime}))^{c},(g^{j}(\bm{X}^{\prime\prime}))^{c}\big{]}_{t}+\sum_{0<s\leq t}\Delta g^{(i)}(\bm{X}^{\prime}_{s})\Delta g^{(j)}(\bm{X}^{\prime\prime}_{s})$
$=\int_{0+}^{t}\Big{(}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})^{\top}\nabla g^{(j)}(\bm{X}^{\prime\prime}_{s-})\Big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{s}$
$\qquad+\sum_{0<s\leq t}\Big{\{}\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}g^{(j)}(\bm{X}^{\prime\prime}_{s})-g^{(j)}(\bm{X}^{\prime\prime}_{s-})\big{)}\Big{\}},$
and
$\big{[}$ $g^{(i)}(\bm{X}^{\prime}_{t}),Y^{(j)}\big{]}_{t}=\big{[}(g^{(i)}(\bm{X}^{\prime}))^{c},(Y^{j})^{c}\big{]}_{t}+\sum_{0<s\leq t}\Delta g^{(i)}(\bm{X}^{\prime}_{s})\Delta Y^{(j)}_{s}$
$=\int_{0+}^{t}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\cdot{\mathrm{d}}[\bm{X}^{\prime},Y^{(j)}]^{c}_{s}+\sum_{0<s\leq t}\Big{\{}\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}Y^{(k)}_{s}-Y^{(k)}_{s-}\big{)}\Big{\}}$.
Consequently, by (2) of Theorem 3.5,
$J_{2}=\overline{{\mathbb{E}}}\Bigg{[}\int_{0+}^{t}$ $\bigg{(}\frac{1}{2}\sum_{i,j=1}^{n}\partial^{2}_{z_{i}z_{j}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{(}\big{(}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\nabla g^{(j)}(\bm{X}^{\prime\prime}_{s-})^{\top}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{s}\Big{)}$ (4.4)
$+\sum_{i=1}^{n}\sum_{k=1}^{d}\partial^{2}_{z_{i}y_{k}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{(}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\cdot{\mathrm{d}}[\bm{X}^{\prime},Y^{(k)}]^{c}_{s}\Big{)}\bigg{)}$
$+\sum_{0<s\leq t}$ $\bigg{(}\frac{1}{2}\sum_{i,j=1}^{n}\partial^{2}_{z_{i}z_{j}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{\{}\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}g^{(j)}(\bm{X}^{\prime}_{s})-g^{(j)}(\bm{X}^{\prime}_{s-})\big{)}\Big{\}}$
$+\sum_{i=1}^{n}\sum_{k=1}^{d}\partial^{2}_{z_{i}y_{k}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{\{}\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}Y^{(k)}_{s}-Y^{(k)}_{s-}\big{)}\Big{\}}\bigg{)}\bigg{|}{\mathcal{F}}\Bigg{]}$.
To deal with $J_{3}$, note that
$\Delta Z^{(i)}_{t}$ $=\big{\langle}\mu_{t},g^{(i)}\big{\rangle}-\lim_{s\nearrow t}\big{\langle}\mu_{s},g^{(i)}\big{\rangle}=\overline{{\mathbb{E}}}\big{[}g^{(i)}(\bm{X}_{t})\big{|}{\mathcal{G}}_{T}\big{]}-\lim_{s\nearrow t}\overline{{\mathbb{E}}}\big{[}g^{(i)}(\bm{X}_{s})\big{|}{\mathcal{G}}_{T}\big{]}$
$=\overline{{\mathbb{E}}}\big{[}g^{(i)}(\bm{X}_{t})\big{|}{\mathcal{G}}_{T}\big{]}-\overline{{\mathbb{E}}}\big{[}\lim_{s\nearrow t}g^{(i)}(\bm{X}_{s})\big{|}{\mathcal{G}}_{T}\big{]}=\overline{{\mathbb{E}}}\big{[}g^{(i)}(\bm{X}_{t})-g^{(i)}(\bm{X}_{t-})\big{|}{\mathcal{G}}_{T}\big{]}$
$=\overline{{\mathbb{E}}}\big{[}g^{(i)}(\bm{X}^{\prime}_{t})-g^{(i)}(\bm{X}^{\prime}_{t-})\big{|}{\mathcal{F}}\big{]},$
where the third equality holds by the conditional dominated convergence theorem (Lemma 2.5). Therefore, we can rewrite $J_{3}$ as
$J_{3}=\sum_{0<s\leq t}\bigg{\{}\overline{{\mathbb{E}}}\bigg{[}$ $\frac{1}{2}\sum_{i,j=1}^{n}\partial^{2}_{z_{i}z_{j}}f(\bm{Y}_{s-},\bm{Z}_{s-})\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}g^{(j)}(\bm{X}^{\prime\prime}_{s})-g^{(j)}(\bm{X}^{\prime\prime}_{s-})\big{)}$ (4.5)
$+\sum_{i=1}^{n}\partial_{z_{i}}f(\bm{Y}_{s-},\bm{Z}_{s-})\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\bigg{|}{\mathcal{F}}\bigg{]}\bigg{\}}$
$+\sum_{0<s\leq t}\overline{{\mathbb{E}}}\bigg{[}$ $\sum_{i=1}^{n}\sum_{k=1}^{d}\partial^{2}_{z_{i}y_{k}}f(\bm{Y}_{s-},\bm{Z}_{s-})\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}Y^{(k)}_{s}-Y^{(k)}_{s-}\big{)}\bigg{|}{\mathcal{F}}\Bigg{]}$,
since $\partial^{2}_{z_{i}z_{j}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $\partial_{z_{i}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $\partial^{2}_{z_{i}y_{k}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $Y^{(k)}_{s}$ and $Y^{(k)}_{s-}$ are $\mathcal{F}$}-measurable random variables for $s\in(0,t]$, $1\leq i,j\leq n$, and $1\leq k\leq d$.
Recall that $\Phi(\mu,\bm{y})=f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)}$, simple calculation of linear derivatives yields a version of the derivatives by
$\nabla_{\bm{y}}\Phi(\mu,\bm{y})=\nabla_{\bm{y}}f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)},\bm{y}\rangle\big{)},\quad\nabla^{2}_{0}\Phi(\mu,\bm{y})=\nabla^{2}_{\bm{y}\bm{y}}f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)},$
$\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})=\sum_{i=1}^{n}\partial_{z_{i}}f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)}g^{(i)}(\bm{x}_{1}),$
$\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})=\sum_{i,j=1}^{n}\partial_{z_{i}z_{j}}f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)}g^{(i)}(\bm{x}_{1})g^{(j)}(\bm{x}_{2}).$
Therefore, we can rewrite $J_{1}$ as
$\;\overline{{\mathbb{E}}}\bigg{[}\int_{0+}^{t}\bigg{(}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s},\bm{X}^{\prime}_{s}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{t}\bigg{)}$ (4.6)
$+\sum_{0<s\leq t}\bigg{(}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{)}\bigg{|}{\mathcal{F}}\bigg{]},$
$J_{2}$ as
$\overline{{\mathbb{E}}}\bigg{[}$ $\int_{0+}^{t}\bigg{(}\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{t}+\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]^{c}_{s}\bigg{)}$ (4.7)
$+\sum_{0<s\leq t}\bigg{(}\frac{1}{2}\bigg{(}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$
$\qquad\qquad\qquad-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}\bigg{)}$
$\qquad\qquad\qquad+\bigg{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{)}\cdot\Delta\bm{Y}_{s}\bigg{)}\bigg{|}{\mathcal{F}}\bigg{]},$
and $J_{3}$ as
$\sum_{0<s\leq t}\bigg{(}\frac{1}{2}\overline{{\mathbb{E}}}$ $\bigg{[}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$ (4.8)
$-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}\bigg{|}{\mathcal{F}}\bigg{]}$
$+$ $\overline{{\mathbb{E}}}\bigg{[}\bigg{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{)}\cdot\Delta\bm{Y}_{s}\bigg{|}{\mathcal{F}}\bigg{]}$
$+$ $\overline{{\mathbb{E}}}\bigg{[}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}{\mathcal{F}}\bigg{]}\bigg{)}.$
Using the argument at the end of [31, Section 3.2.1], one can cancel out redundant terms in $J_{2}$ and $J_{3}$ to get the first summation in the right hand side of (2.3). Therefore, combining (4.6), (4.7), and (4.8) and plugging them back into (4.2) completes the proof.
4.2. Step 2: localization argument.
Take a sequence of $\mathbb{F}$}-stopping time $\{\tau_{n}\}_{n\in{\mathbb{N}}}$ such that $\tau_{n}\to T$, in order to prove (2.3) for $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ with the process $(\mu,\bm{Y})=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$, it suffices to prove (2.3) for $\Phi$ with the truncated process $(\mu^{(\tau_{n})},\bm{Y}^{(\tau_{n})}):=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$. Indeed, suppose (2.3) holds for the truncated process, for $t<T$. Let $n\to\infty$, it is not hard to see that the left hand side of (2.3) converges and the last three terms of the right hand side of (2.3) converge to the desired equation.
Now, let us split the conditional expectation term of the right hand side of (2.3) into several terms. First, in order to prove that
$\overline{{\mathbb{E}}}\bigg{[}\int_{0+}^{t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\bigg{|}{\mathcal{F}}\bigg{]}\to\overline{{\mathbb{E}}}\bigg{[}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$ (4.9)
a.s. as $n\to\infty$, we separate it by,
$\int_{0+}^{t\wedge\tau_{n}}$ $\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}=$ (4.10)
$\int_{0+}^{t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{s}-\sum_{0<s\leq t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot\Delta\bm{X}^{\prime}_{s}.$
Note that the first term is uniformly bounded by
$\sup_{0\leq s\leq T}\bigg{|}\int_{0+}^{s}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{u}\bigg{|},$
and
$\overline{{\mathbb{E}}}\bigg{[}\sup_{0\leq s\leq T}\bigg{|}\int_{0+}^{s}$ $\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{u}\bigg{|}\bigg{]}=\bigg{\|}\int_{0+}^{\cdot}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{u}\bigg{\|}_{{\mathcal{S}}_{1}}$
$\leq c_{1}\bigg{\|}\int_{0+}^{\cdot}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{u}\bigg{\|}_{{\mathcal{H}}_{1}}$
$\leq c_{1}\bigg{\|}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{\cdot},\bm{Y}_{\cdot},\bm{X}^{\prime}_{\cdot}\big{)}\bigg{\|}_{{\mathcal{S}}_{p^{\prime}}}\big{\|}\bm{X}^{\prime}\big{\|}_{{\mathcal{H}}_{p}}\leq c_{1}C\big{\|}|\bm{X}|^{p-1}+|\bm{Y}|^{p-1}+1\big{\|}_{{\mathcal{S}}_{p^{\prime}}}\big{\|}\bm{X}\big{\|}_{{\mathcal{H}}_{p}}$
$\leq c_{p}c_{1}C\big{(}\big{\|}\bm{X}\big{\|}^{p}_{{\mathcal{H}}_{p}}+\big{\|}\bm{Y}\big{\|}^{p}_{{\mathcal{H}}_{p}}+1\big{)}\big{\|}\bm{X}\big{\|}_{{\mathcal{H}}_{p}}<\infty,$
where $p^{\prime}=\frac{p}{p-1}$, $c_{1}$ and $c_{p}$ are constants in Proposition 2.2, and $C$ is a generic constant that may vary line by line. The first inequality is due to Proposition 2.2, the second inequality is by Proposition 2.3, the third inequality is from $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, and the last one holds because $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. Meanwhile, the second term of the right hand side of (4.10) is bounded by
$\sup_{0\leq s\leq T}\bigg{|}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},$ $\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}\cdot\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|\leq C\sup_{0\leq s\leq T}\big{(}1+|\bm{X}_{s}|^{p-1}+|\bm{Y}_{s}|^{p-1}\big{)}\cdot\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|$ (4.11)
$\leq C\bigg{(}1+\sup_{0\leq s\leq T}|\bm{X}_{s}|^{p}+\sup_{0\leq s\leq T}|\bm{Y}_{s}|^{p}+\bigg{(}\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|\bigg{)}^{p}\bigg{)},$
where $C$ is a generic constant, and the right hand side of the above inequality is integrable because $\|\bm{X}\|_{{\mathcal{S}}_{p}}\leq c_{p}\|\bm{X}\|_{{\mathcal{H}}_{p}}<\infty$, $\|\bm{Y}\|_{{\mathcal{S}}_{p}}\leq c_{p}\|\bm{Y}\|_{{\mathcal{H}}_{p}}<\infty$, and by the assumption of $\bm{X}$. Therefore, by the conditional dominated convergence theorem (see Lemma 2.5), we conclude (4.9).
Next, to prove that
$\overline{{\mathbb{E}}}\bigg{[}\int_{0+}^{t\wedge\tau_{n}}$ $\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}\bigg{|}{\mathcal{F}}\bigg{]}$ (4.12)
$\to\overline{{\mathbb{E}}}\bigg{[}\int_{0+}^{t}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
a.s. as $n\to\infty$, note that
$\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}$
is bounded by
$\sup_{0\leq s\leq T}\bigg{|}\int_{0+}^{s}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{u}\bigg{|}\leq\sup_{0\leq s\leq T}\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}\cdot[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T},$
and
$\overline{{\mathbb{E}}}\bigg{[}\sup_{0\leq s\leq T}\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}$ $\bigg{|}\cdot[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T}\bigg{]}\leq\bigg{\|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{\cdot},\bm{Y}_{\cdot},\bm{X}^{\prime}_{\cdot}\big{)}\bigg{\|}_{{\mathcal{S}}_{p^{\prime\prime}}}\big{\|}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T}\big{\|}_{L_{p/2}}$
$\leq C\Big{\|}1+|\bm{X}^{\prime}|^{p-2}+|\bm{Y}|^{p-2}\Big{\|}_{{\mathcal{S}}_{p^{\prime\prime}}}\big{\|}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T}\big{\|}_{L_{p/2}}$
$\leq c_{p^{\prime\prime}}C\Big{(}1+\|\bm{X}^{\prime}\|^{p}_{{\mathcal{H}}_{p}}+\|\bm{Y}\|^{p}_{{\mathcal{H}}_{p}}\Big{)}\big{\|}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T}\big{\|}_{L_{p/2}}<\infty,$
where $c_{p^{\prime\prime}}$ is the constant in Proposition 2.2, $C$ is a generic constant that may vary from line to line and $p^{\prime\prime}=\frac{p-2}{2p}$. Thus, (4.12) holds by Lemma 2.5. Similar arguments holds also for the convergence of $\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]_{s}$ and $\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]_{s}$.
Next, let us deal with the convergence of
$\sum_{0<s\leq t\wedge\tau_{n}}\bigg{(}$ $\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$ (4.13)
$-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}\bigg{)}\mathbbm{1}_{\{\mu_{s}=\mu_{s-}\}}.$
Applying the fundamental theorem of calculus for the linear derivative, we have
$\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$
$-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}$
$=\int_{0}^{1}$ $\int_{0}^{1}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}+\lambda(\bm{X}^{\prime}_{s}-\bm{X}^{\prime}_{s-}),\bm{X}^{\prime\prime}_{s-}+\gamma(\bm{X}^{\prime\prime}_{s}-\bm{X}^{\prime\prime}_{s-})\big{)}:\Big{(}\big{(}\Delta\bm{X}^{\prime}_{s}\big{)}^{\top}\Delta\bm{X}^{\prime\prime}_{s}\Big{)}{\mathrm{d}}\lambda{\mathrm{d}}\gamma,$
and the right hand side of the above equation is bounded by
$C\cdot|\Delta\bm{X}^{\prime}_{s}|\cdot|\Delta\bm{X}^{\prime\prime}_{s}|\cdot\Big{(}1+|\bm{X}^{\prime}_{s}|^{p-2}+|\bm{X}^{\prime\prime}_{s}|^{p-2}\Big{)}.$
Therefore, (4.13) is bounded by
$\sup_{0\leq s\leq T}\Big{(}1+|\bm{X}^{\prime}_{s}|^{p-2}+|\bm{X}^{\prime\prime}_{s}|^{p-2}\Big{)}\big{(}\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|\cdot|\Delta\bm{X}^{\prime\prime}_{s}|\big{)}$
$\leq C\bigg{(}\sup_{0\leq t\leq T}\Big{(}1+|\bm{X}^{\prime}_{t}|^{p-2}+|\bm{X}^{\prime\prime}_{t}|^{p-2}\Big{)}^{\frac{p}{p-2}}+\Big{(}\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|^{2}+\sum_{0<s\leq T}|\Delta\bm{X}^{\prime\prime}_{s}|^{2}\Big{)}^{\frac{p}{2}}\bigg{)}$
$\leq C\bigg{(}\sup_{0\leq t\leq T}\Big{(}1+|\bm{X}^{\prime}_{t}|^{p}+|\bm{X}^{\prime\prime}_{t}|^{p}\Big{)}+\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|^{p}+\sum_{0<s\leq T}|\Delta\bm{X}^{\prime\prime}_{s}|^{p}\bigg{)},$
where $C$ is a generic constant. The right hand side of the above inequality is integrable and the convergence of the conditional expectation of (4.13) is ensured by Lemma 2.5. Similar arguments can be applied to deduce the convergence of the conditional expectation of $\big{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\big{)}\cdot\Delta\bm{Y}_{s}$ and $\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}$, and thus we conclude that (2.3) holds for the original process $(\mu,\bm{Y})$ for $0\leq t<T$.
For the case of $t=T$, taking $n\to\infty$, then the left hand side equals $\Phi(\mu_{T-},\bm{Y}_{T-})-\Phi(\mu_{0},\bm{Y}_{0})$ while all terms on the right hand side are evaluated at $T$ except for the last term which is
$\sum_{0<s<t}\Big{(}\Phi(\mu_{s},\bm{Y}_{s})-\Phi(\mu_{s-},\bm{Y}_{s-})\Big{)},$
adding $\Phi(\mu_{T},\bm{Y}_{T})-\Phi(\mu_{T-},\bm{Y}_{T-})$ to both sides, we finish the proof of the claim.
4.3. Step 3: cylindrical functions are dense in $C^{2,2}({\mathbb{R}}^{d}\times\mathcal{P}_{p}({\mathbb{R}}^{d}))$.
The following result slightly generalizes that in [21, Theorem 4.4, Theorem 4.10, Section 4] as it also includes the convergences of the associated linear derivatives.
Proposition 4.2.
Let $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, then there exists a sequence of $C^{2,2}$ cylinder functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$ such that one has the point-wise convergence
$\bigg{(}f_{n},\nabla_{\bm{y}}f_{n}$ $,\nabla_{\bm{y}}^{2}f_{n},\frac{\delta f_{n}}{\delta\mu},\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu},\nabla^{2}_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu},\nabla_{\bm{x}_{1}\bm{y}}\frac{\delta f_{n}}{\delta\mu},\nabla_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}f_{n}}{(\delta\mu)^{2}}\bigg{)}$ (4.14)
$\to\bigg{(}\Phi,\nabla_{\bm{y}}\Phi,\nabla_{\bm{y}}^{2}\Phi,\frac{\delta\Phi}{\delta\mu},\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu},\nabla^{2}_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu},\nabla_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu},\nabla_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\bigg{)}$
as $n\to\infty$, for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$. Moreover, there exists a constant $c>0$ such that
$|\nabla_{\bm{y}}f_{n}|(\mu,\bm{y})\leq c(1+|\bm{y}|^{p-1}),\qquad|\nabla^{2}_{\bm{y}}f_{n}|(\mu,\bm{y})$ $\leq c(1+|\bm{y}|^{p-2}),$ (4.15)
$\bigg{|}\frac{\delta f_{n}}{\delta\mu}\bigg{|}\leq c\big{(}1+|\bm{x}_{1}|^{p}+|\bm{y}|^{p}\big{)},\qquad\bigg{|}\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-1}+|\bm{y}|^{p-1}\big{)},$
$\bigg{|}\nabla^{2}_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{y}|^{p-2}\big{)},\qquad\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{{p-2}}+|\bm{y}|^{p-2}\big{)},$
$\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}f_{n}}{(\delta\mu)^{2}}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{x}_{2}|^{p-2}+|\bm{y}|^{p-2}\big{)},$
for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$.
The proof is similar to that of [21, Section 4]. Following their construction procedure, fix $n\in{\mathbb{N}}$ and cover the ball $B_{n}:=\{\bm{x}\in{\mathbb{R}}^{d}:|\bm{x}|\leq n\}$ by finitely many open sets of diameter at most $1/n$, denoted by $U^{n}_{i}$, with $i=1,...,N_{n}$. Set $U^{n}_{0}={\mathbb{R}}^{d}\setminus B_{n}$ to get an open cover in ${\mathbb{R}}^{d}$. Then, fix points ${\bm{x}}^{n}_{i}$ in $\overline{U^{n}_{i}}$ with the minimal norm for $i=0,1,..,N_{n}$. Next, define a family of functions $\{\psi^{n}_{i}\}$ as a partition of unity subordinate to $\{U^{n}_{i}\}$. (That is, each $\psi^{n}_{i}$ is a continuous function supported on $U^{n}_{i}$, such that $\sum_{i=0}^{N_{n}}\psi^{n}_{i}(\bm{x})=1$ for all $\bm{x}\in{\mathbb{R}}^{d}$). Finally, define the operator $T_{n}:C_{b}({\mathbb{R}}^{d})\to C_{b}({\mathbb{R}}^{d})$ such that for any function $\varphi$ on ${\mathbb{R}}^{d}$,
$T_{n}\varphi(x)=\sum_{i=0}^{N_{n}}\varphi({\bm{x}}^{n}_{i})\psi^{n}_{i}(\bm{x}).$
Its adjoint operator $T_{n}^{*}:{\mathcal{P}}_{p}({\mathbb{R}}^{d})\to{\mathcal{P}}_{p}({\mathbb{R}}^{d})$ and the tensor square operator $T_{n}^{\otimes 2}:C_{b}({\mathbb{R}}^{2d})\to C_{b}({\mathbb{R}}^{2d})$ are defined accordingly. Now, define $f_{n}(\mu,\bm{y}):=\Phi(T_{n}^{*}\mu,\bm{y})$, then $f_{n}$ are $C^{2,2}$ cylindrical functions and the key identities to be established are
$\frac{\delta f_{n}}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})=T_{n}\Big{(}\frac{\delta\Phi}{\delta\mu}(T_{n}^{*}\mu,\bm{y},\cdot)\Big{)}(\bm{x}_{1}),\quad\frac{\delta^{2}f_{n}}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})=T_{n}^{\otimes 2}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(T_{n}^{*}\mu,\bm{y},\cdot,\cdot)(\bm{x}_{1},\bm{x}_{2}),$
for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$ and $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$ (see [21, Lemma 4.6, Lemma 4.12]). To this end, note that the construction of $T_{n}$ ensures
$T_{n}\nabla_{\bm{x}_{1}}\varphi(\bm{x}_{1})=\nabla_{\bm{x}_{1}}T_{n}\varphi(\bm{x}_{1}),\quad T_{n}^{\otimes 2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\psi(\bm{x}_{1},\bm{x}_{2})=\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}T_{n}\psi(\bm{x}_{1},\bm{x}_{2}),$
for all $\varphi\in C^{1}_{b}({\mathbb{R}}^{d})$ and $\psi\in C^{2}_{b}({\mathbb{R}}^{2d}).$ Thus repeating the proof of [21, Theorem 4.4, Theorem 4.10] finishes the proof of the proposition.
4.4. Final step
Define $\tau_{k}:=\inf\{t\in[0,T]:|\bm{Y}_{t}|>k\}$, then by Step 1, Itô’s formula (2.3) holds for the truncated process $(\mu^{(\tau_{n})},\bm{Y}^{(\tau_{n})}):=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$ with the sequences of cylindrical functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$. By Step 2, the localization argument shows that Itô’s formula (2.3) holds for the sequences of cylindrical functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$.
In order to pass the limit from $f_{n}$ to $\Phi$, let us consider the first term in the right hand side of (2.3) and the remaining terms follow similarly. For the first term, we need to prove that there exists a subsequence $\{n_{k}\}_{k\in{\mathbb{N}}^{+}}$ such that
${\mathbb{E}}\bigg{[}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta f_{n_{k}}}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\Big{|}{\mathcal{F}}\bigg{]}\to{\mathbb{E}}\bigg{[}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\Big{|}{\mathcal{F}}\bigg{]},$ (4.16)
as $k\to\infty.$ To this end, we first prove that
$\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\overset{L^{1}}{\to}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s},$ (4.17)
as $n\to\infty$. In fact, by Proposition 4.2, we have
$\bigg{|}|\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}\leq c\big{(}1+|\bm{X}_{s-}|^{p-1}+|\bm{Y}_{s-}|^{p-1}\big{)},$
where the ${\mathcal{S}}_{p^{\prime}}$ norm of the right hand side is finite with $p^{\prime}=\frac{p}{p-1}$. Now (4.17) follows from $\|\bm{X}\|_{{\mathcal{H}}^{p}}<\infty$ and Lemma 2.5.
Next, since
${\mathbb{E}}\bigg{[}\Big{|}$ ${\mathbb{E}}\Big{[}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\Big{|}{\mathcal{F}}\Big{]}-{\mathbb{E}}\Big{[}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\Big{|}{\mathcal{F}}\Big{]}\Big{|}\bigg{]}$
$\leq{\mathbb{E}}\bigg{[}\Big{|}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}-\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\Big{|}\bigg{]}\to 0,$
and $L_{1}$ convergences implies convergence in subsequence, (4.16) holds. The remaining terms follow similarly due to the uniform bound (4.15), and thus the proof of Theorem 2.11 is complete.
5. Proof of Results in Section 3
5.1. Proof of Theorem 3.1
To prove the theorem, we first provide a constructive proof for the existence of conditionally independent copy claimed in Theorem 3.1. For ease of exposition, we will show how to enlarge the probability space with two independent copies, and the case with a finite number of copies can be easily adapted.
Recall the discussions in Section 3, and consider the probability space $\overline{\Omega}$, the $\sigma$-algebra $\overline{{\mathcal{F}}}$ defined by
$\overline{\Omega}=\Omega^{3}=\{(\omega_{0},\omega_{1},\omega_{2})|\omega_{0},\omega_{1},\omega_{2}\in\Omega\},\qquad\overline{{\mathcal{F}}}=\sigma\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}\},$ (5.1)
and a probability measure $\overline{{\mathbb{P}}}$ defined as follows: for $A_{0},A_{1},A_{2}\in{\mathcal{F}}$,
$\overline{{\mathbb{P}}}\big{(}A_{0}\times A_{1}\times A_{2}\big{)}:={\mathbb{E}}\Big{[}\mathbbm{1}_{A_{0}}{\mathbb{P}}(A_{1}|{\mathcal{G}}){\mathbb{P}}(A_{2}|{\mathcal{G}})\Big{]}.$ (5.2)
Since the set $\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}\}$ is a $\pi$-system, the measure on $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ is well defined.
Next, take $\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$ for the natural extension in this enlarged probability space as
$\widetilde{\mathcal{F}}:=\{A\times\Omega\times\Omega:A\in{\mathcal{F}}\},\quad\widetilde{\mathcal{G}}:=\{A\times\Omega\times\Omega:A\in{\mathcal{G}}\},$ (5.3)
and define $\mathcal{F}$$^{\prime}$, $\mathcal{F}$$^{\prime\prime}$ as
${\mathcal{F}}^{\prime}:=\{\Omega\times A\times\Omega:A\in{\mathcal{F}}\},\quad{\mathcal{F}}^{\prime\prime}:=\{\Omega\times\Omega\times A:A\in{\mathcal{F}}\}.$ (5.4)
In this framework, we have the following proposition.
Proposition 5.1.
Consider $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ defined by (5.1) and (5.2) and the sub-algebras $\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$, $\mathcal{F}$$^{\prime}$, and $\mathcal{F}$$^{\prime\prime}$ defined by (5.3) and (5.4). Suppose $A,B\in{\mathcal{F}}$, and let $\widetilde{A}=A\times\Omega\times\Omega$, $A^{\prime}=\Omega\times A\times\Omega$, $A^{\prime\prime}=\Omega\times\Omega\times A$, $B^{\prime}=\Omega\times B\times\Omega$, and $C^{\prime\prime}=\Omega\times\Omega\times C$. Then
  1. (1)
    For $\overline{\omega}=(\omega_{0},\omega_{1},\omega_{2})\in\overline{\Omega}$,
[…]
  2. (2)
    Given $\widetilde{\mathcal{G}}$, $\widetilde{A},A^{\prime},$ and $A^{\prime\prime}$ have the same conditional law. That is,
[…]
  3. (3)
    Given $\widetilde{\mathcal{G}}$, $\widetilde{\mathcal{F}}$$\mathcal{F}$$^{\prime}$, and $\mathcal{F}$$^{\prime\prime}$ are conditionally independent. That is,
    $\overline{{\mathbb{P}}}\big{(}\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)}=\overline{{\mathbb{P}}}\big{(}\widetilde{A}|\widetilde{\mathcal{G}}\big{)}\overline{{\mathbb{P}}}\big{(}B^{\prime}|\widetilde{\mathcal{G}}\big{)}\overline{{\mathbb{P}}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)},\text{ a.s. under }\overline{{\mathbb{P}}}.$
    Moreover, given $\widetilde{\mathcal{F}}$, then $\mathcal{F}$$^{\prime}$ and $\mathcal{F}$$^{\prime\prime}$ are conditionally independent. That is,
    $\overline{{\mathbb{P}}}\big{(}B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)}=\overline{{\mathbb{P}}}\big{(}B^{\prime}|\widetilde{\mathcal{F}}\big{)}\overline{{\mathbb{P}}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)},\text{ a.s. under }\overline{{\mathbb{P}}}.$
Proof.
(of Proposition 5.1). To prove (1), note that ${\mathbb{P}}[A|{\mathcal{G}}]$ is $\mathcal{G}$} measurable in $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$, then ${\mathbb{P}}[A|{\mathcal{G}}](\omega_{1})$ as a random variable in $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}},\overline{{\mathbb{P}}})$ is $\widetilde{\mathcal{G}}-$measurable. For any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$. By the definition of $\overline{{\mathbb{P}}}$ in ((5.2)) and the property of conditional expectation,
$\overline{{\mathbb{P}}}\big{[}\widetilde{A}\cap\widetilde{G}\big{]}=\overline{{\mathbb{P}}}\big{[}(A\cap G)\times\Omega\times\Omega\big{]}={\mathbb{P}}[A\cap G]={\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{G}\Big{]}=\int_{\Omega}{\mathbb{P}}[A|{\mathcal{G}}](\omega)\mathbbm{1}_{G}(\omega){\mathbb{P}}({\mathrm{d}}\omega).$
Again by (5.2), $\overline{{\mathbb{P}}}({\mathrm{d}}\omega_{1}\times\Omega\times\Omega)={\mathbb{P}}({\mathrm{d}}\omega_{1})$, therefore,
$\overline{{\mathbb{P}}}\big{[}\widetilde{A}\cap\widetilde{G}\big{]}$ $=\int_{\Omega^{3}}{\mathbb{P}}[A|{\mathcal{G}}](\omega_{0})\mathbbm{1}_{G}(\omega_{0})\overline{{\mathbb{P}}}({\mathrm{d}}\omega_{0}{\mathrm{d}}\omega_{1}{\mathrm{d}}\omega_{2})$ (5.5)
$=\int_{\Omega^{3}}{\mathbb{P}}[A|{\mathcal{G}}](\omega_{0})\mathbbm{1}_{\widetilde{G}}(\omega_{0},\omega_{1},\omega_{2})\overline{{\mathbb{P}}}({\mathrm{d}}\omega_{0}{\mathrm{d}}\omega_{1}{\mathrm{d}}\omega_{2})=\overline{{\mathbb{E}}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]}$,
which leads to ${\mathbb{P}}[A|{\mathcal{G}}](\omega_{0})=\overline{{\mathbb{P}}}[\widetilde{A}|\widetilde{\mathcal{G}}](\omega_{0},\omega_{1},\omega_{2})$ a.s. under $\overline{P}$.
To prove (2), for any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$. Then,
$\overline{{\mathbb{P}}}\big{[}A^{\prime}\cap\widetilde{G}\big{]}=\overline{{\mathbb{P}}}\big{[}G\times A\times\Omega\big{]}={\mathbb{E}}\big{[}\mathbbm{1}_{G}{\mathbb{P}}(A|{\mathcal{G}})\big{]}=\overline{{\mathbb{E}}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{P}}}[\widetilde{A}|\widetilde{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]}$,
[…]
$\overline{{\mathbb{P}}}(A^{\prime}\cap\widetilde{F})=\overline{{\mathbb{P}}}\big{[}F\times A\times\Omega\big{]}={\mathbb{E}}\big{[}\mathbbm{1}_{F}{\mathbb{P}}(A|{\mathcal{G}})\big{]}=\overline{{\mathbb{E}}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{F}}\Big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{P}}}[\widetilde{A}|\widetilde{\mathcal{G}}]\mathbbm{1}_{\widetilde{F}}\Big{]}$,
[…]
To prove (3), for any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$, and
[…]
$={\mathbb{E}}\Big{[}{\mathbb{E}}[\mathbbm{1}_{A\cap G}|{\mathcal{G}}]{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}={\mathbb{E}}\Big{[}\mathbbm{1}_{G}{\mathbb{P}}(A|{\mathcal{G}}){\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}.$
[…]
Therefore,
$\overline{{\mathbb{P}}}\big{(}\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)}=\overline{{\mathbb{P}}}(\widetilde{A}|\widetilde{\mathcal{G}})\overline{{\mathbb{P}}}(B^{\prime}|\widetilde{\mathcal{G}})\overline{{\mathbb{P}}}(C^{\prime\prime}|\widetilde{\mathcal{G}}),\text{ a.s. under }\overline{{\mathbb{P}}}.$
As for the second statement of (3), for any $\widetilde{F}\in\widetilde{\mathcal{F}}$, there exists an $F\in{\mathcal{F}}$ such that $\widetilde{F}=F\times\Omega\times\Omega\in\widetilde{\mathcal{F}}$, and by (5.2),
[…]
$=\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{\widetilde{F}}{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\big{]}=\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{\widetilde{F}}\overline{{\mathbb{P}}}(B^{\prime}|\widetilde{\mathcal{F}})\overline{{\mathbb{P}}}(C^{\prime\prime}|\overline{{\mathcal{F}}})\big{]},$
where the the third equality follows from the same calculation as (5.5) and the last equation holds by (1) and (2). Therefore,
$\overline{{\mathbb{P}}}\big{(}B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)}=\overline{{\mathbb{P}}}\big{(}B^{\prime}|\widetilde{\mathcal{F}}\big{)}\overline{{\mathbb{P}}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)},\text{ a.s. under }\overline{{\mathbb{P}}}.$
Proof.
(of Theorem 3.1). For any random variable $R$ taking values in a standard Borel space $(E,\mathcal{E})$, recall that the natural extension $\widetilde{R}$ and the two conditionally independent copies $R^{\prime}$ and $R^{\prime\prime}$ of $R$ are defined by
$\widetilde{R}(\omega_{0},\omega_{1},\omega_{2})=R(\omega_{0}),\;R^{\prime}(\omega_{0},\omega_{1},\omega_{2})=R(\omega_{1}),\;R^{\prime\prime}(\omega_{0},\omega_{1},\omega_{2})=R(\omega_{2}).$
Then
[…]
Remark 5.2.
Our construction is inspired by the proof of Yamada-Watanabe theorem (see for instance [39, 42]) and the theory of weak and strong solutions of stochastic differential equations (see for instance [37, 43]). Note that in these existing approaches, the common information $\mathcal{G}$} is modeled as part of the stochastic input. Moreover, in the context of weak and strong solutions of SDEs, one often uses the existence of regular conditional distributions to construct an extended probability space by embedding different weak solutions into a common space. To illustrate this, take the example of $(X^{(1)},W^{(1)})$ and $(X^{(2)},W^{(2)})$, let $\nu({\mathrm{d}}w)$ denote the common law of a Brownian motion, and define the regular conditional law $\eta^{(i)}({\mathrm{d}}x;w):=\text{Law}(X^{(i)}|W^{(i)}=w)$ for $i=1,2$. In such a setting, an extended probability measure is often defined (informally) by ${\mathbb{P}}({\mathrm{d}}x^{(1)}{\mathrm{d}}x^{(2)}{\mathrm{d}}w):=\eta^{(1)}({\mathrm{d}}x^{(1)};w)\cdot\eta^{(2)}({\mathrm{d}}x^{(2)};w)\nu({\mathrm{d}}w).$
Motivated by these constructions, we consider a more general setting in which the common information $\mathcal{G}$} may be abstract and regular conditional distributions may not be available. We define conditionally independent copies of some random variable $X$ directly by specifying the measure on the extended space supporting $X,X^{\prime}$ such that
${\mathbb{P}}\big{(}(X,X^{\prime})\in A\times B\big{)}:={\mathbb{E}}[\mathbbm{1}_{X\in A}\cdot{\mathbb{P}}(X\in B|{\mathcal{G}})],$
which remains consistent with the spirit of the classical constructions while avoiding the explicit use of regular conditional distributions.
Proof of Corollary 3.2. It suffices to show that the extension of the sub-filtration $\widetilde{\mathbb{G}}\subset\widetilde{\mathbb{F}}$ satisfies
${\widetilde{\mathcal{F}}}_{t}\perp\!\!\!\perp{\widetilde{\mathcal{G}}}_{T}\big{|}{\widetilde{\mathcal{G}}}_{t},\text{ for all }0\leq t\leq T.$
For any $\widetilde{F}_{t}\in\widetilde{\mathcal{F}}_{t}$, $\widetilde{G}_{t}\in{\widetilde{\mathcal{G}}}_{t}$ and $\widetilde{G}_{T}\in{\widetilde{\mathcal{G}}}_{T}$, there exist $F_{t}\in{\mathcal{F}}_{t}$, $G_{t}\in{\mathcal{G}}_{t}$, and $G_{T}\in{\mathcal{G}}_{T}$ such that $\widetilde{F}_{t}=F_{t}\times\Omega\times\Omega$, $\widetilde{G}_{t}=G_{t}\times\Omega\times\Omega$ and $\widetilde{G}_{T}=G_{T}\times\Omega\times\Omega$. Then,
$\overline{{\mathbb{P}}}(\widetilde{F}_{t}\cap\widetilde{G}_{t}\cap\widetilde{G}_{T})$ $=\overline{{\mathbb{P}}}((F_{t}\cap G_{t}\cap G_{T})\times\Omega\times\Omega)={\mathbb{P}}(F_{t}\cap G_{t}\cap G_{T})$
$={\mathbb{E}}\big{[}{\mathbb{P}}(F_{t}\cap G_{T}|{\mathcal{G}}_{t})\mathbbm{1}_{G_{t}}\big{]}={\mathbb{E}}\big{[}{\mathbb{P}}(F_{t}|{\mathcal{G}}_{t}){\mathbb{P}}(G_{T}|{\mathcal{G}}_{t})\mathbbm{1}_{G_{t}}\big{]}$
$=\overline{{\mathbb{E}}}\big{[}{\mathbb{P}}(F_{t}|{\mathcal{G}}_{t}){\mathbb{P}}(G_{T}|{\mathcal{G}}_{t})\mathbbm{1}_{\widetilde{G}_{t}}\big{]}=\overline{{\mathbb{E}}}\big{[}\overline{{\mathbb{P}}}(\widetilde{F}_{t}|\widetilde{\mathcal{G}}_{t})\overline{{\mathbb{P}}}(\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t})\mathbbm{1}_{\widetilde{G}_{t}}\big{]},$
where the fourth equality holds by ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$, the fifth equality follows from the same calculation as (5.5), and the last equality is by (2) in Proposition 5.1. Therefore,
$\overline{{\mathbb{P}}}(\widetilde{F}_{t}\cap\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t})=\overline{{\mathbb{P}}}(\widetilde{F}_{t}|\widetilde{\mathcal{G}}_{t})\overline{{\mathbb{P}}}(\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t}).$
5.2. Proof of Proposition 3.3
For (1), let us assume that ([51, Theorem 1, Chapter 4])
$X_{t}=M_{t}+V_{t},$
where $M_{t}$ is an $\mathbb{F}$}-local martingale and $V_{t}$ is a finite variation process. To prove that $X$ is an $\overline{{\mathbb{F}}}$-semimartingale, it suffices to show that $M_{t}$ is an $\overline{{\mathbb{F}}}$-local martingale. Let $\{\tau_{k}\}_{k\in{\mathbb{N}}^{+}}$ be a sequence of $\mathbb{F}$}-stopping times such that $M^{\tau_{k}}:=(M_{t\wedge\tau_{k}})_{t\in[0,T]}$ is an $\mathbb{F}$}-martingale for all $k\in{\mathbb{N}}^{+}$, and $\tau_{k}\to\infty$ as $k\to\infty$. We claim that
$\overline{{\mathbb{E}}}[M_{t\wedge\tau_{k}}|\overline{{\mathcal{F}}}_{s}]=M_{s\wedge\tau_{k}}.$
Recall that
$\overline{{\mathcal{F}}}_{s}=\sigma\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}_{s}\},$
it suffices to show that
$\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}=\overline{{\mathbb{E}}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}$ (5.6)
for all $A_{0},A_{1},A_{2}\in{\mathcal{F}}_{s}$. Note that $M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}$, $\mathbbm{1}_{\Omega\times A_{1}\times\Omega}$ and $\mathbbm{1}_{\Omega\times\Omega\times A_{2}}$ are conditionally independent given ${\mathcal{G}}_{T}$ by (3) in Proposition 5.1, we have
$\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{\Omega\times A_{1}\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{\Omega\times\Omega\times A_{2}}\big{|}{\mathcal{G}}_{T}\big{]}\Big{]}.$ (5.7)
By (2) in Proposition 5.1 and ${\mathcal{F}}_{s}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{s}$,
$\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{\Omega\times A_{1}\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{{\mathbb{P}}}(A_{1}|{\mathcal{G}}_{T})=\overline{{\mathbb{P}}}(A_{1}|{\mathcal{G}}_{s}),\text{ and similarly }\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{\Omega\times\Omega\times A_{2}}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{{\mathbb{P}}}(A_{2}|{\mathcal{G}}_{s}).$ (5.8)
Therefore (5.7) becomes
$\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}$ $=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{P}}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{{\mathbb{P}}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]}$ (5.9)
$=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{{\mathbb{P}}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{{\mathbb{P}}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]},$
where the second equality holds by the tower rule. Since $M^{\tau_{k}}$ is an $\mathbb{F}$}-martingale,
$\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}=\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}{\mathbb{E}}[M_{t\wedge\tau_{k}}|{\mathcal{F}}_{s}]\big{|}{\mathcal{G}}_{s}\big{]}=\overline{{\mathbb{E}}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}.$
Therefore,
$\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}$ $=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{{\mathbb{P}}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{{\mathbb{P}}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]}.$
Repeating the calculation as in (5.7), (5.8), and (5.9) yields
$\overline{{\mathbb{E}}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}$ $=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{{\mathbb{P}}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{{\mathbb{P}}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]},$
which leads to (5.6).
To show that $X^{\prime}$ is an $\overline{{\mathbb{F}}}$-semimartingale, by the construction of the extended probability space, we can define $M^{\prime},V^{\prime}$, and $\tau_{k}^{\prime}$ to be respectively the conditionally independent copy of $M,V$, and $\tau_{k}$. Then it suffices to show that
${\mathbb{E}}\big{[}M^{\prime}_{t\wedge\tau^{\prime}_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}={\mathbb{E}}\big{[}M^{\prime}_{s\wedge\tau^{\prime}_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}.$
Note that
$\overline{{\mathbb{E}}}\big{[}M^{\prime}_{t\wedge\tau_{k}^{\prime}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M^{\prime}_{t\wedge\tau^{\prime}_{k}}\mathbbm{1}_{\Omega\times A_{2}\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{P}}}(A_{0}|{\mathcal{G}}_{s})\overline{{\mathbb{P}}}(A_{2}|{\mathcal{G}}_{s})\Big{]}$
$=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{1}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{P}}}(A_{0}|{\mathcal{G}}_{s})\overline{{\mathbb{P}}}(A_{2}|{\mathcal{G}}_{s})\Big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{1}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{{\mathbb{P}}}(A_{0}|{\mathcal{G}}_{s})\overline{{\mathbb{P}}}(A_{2}|{\mathcal{G}}_{s})\Big{]}$
$=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{1}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{{\mathbb{P}}}(A_{0}|{\mathcal{G}}_{s})\overline{{\mathbb{P}}}(A_{2}|{\mathcal{G}}_{s})\Big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{1}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{P}}}(A_{0}|{\mathcal{G}}_{s})\overline{{\mathbb{P}}}(A_{2}|{\mathcal{G}}_{s})\Big{]}$
$=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M^{\prime}_{s\wedge\tau^{\prime}_{k}}\mathbbm{1}_{\Omega\times A_{2}\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{P}}}(A_{0}|{\mathcal{G}}_{s})\overline{{\mathbb{P}}}(A_{2}|{\mathcal{G}}_{s})\Big{]}={\mathbb{E}}\big{[}M^{\prime}_{s\wedge\tau^{\prime}_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]},$
where the first and last equalities mimic the argument in (5.7), (5.8) and (5.9), the second and the sixth equalities hold from (2) in Proposition 5.1, the third and the fifth equalities follow from the tower rule, and the fourth equality is the martingale property. This completes the proof of (1).
For (2a), it suffices to prove that
$\overline{{\mathbb{E}}}\bigg{[}\bigg{(}\sup_{\sigma_{m}}\sum_{i=1}^{k_{m}}\big{|}X^{\mathcal{G}}_{t_{i+1}^{m}}-X^{\mathcal{G}}_{t_{i}^{m}}\big{|}\bigg{)}^{p}\bigg{]}<\infty,$
where the supreme is taken over all partition $\sigma_{m}:\{0=t_{0}^{m}\leq t_{1}^{m}\leq\cdots\leq t_{k_{m}}^{m}=T\}$ of $[0,T]$. Since
$\sum_{i=1}^{k_{m}}\big{|}X^{\mathcal{G}}_{t_{i+1}^{m}}-X^{\mathcal{G}}_{t_{i}^{m}}\big{|}=\sum_{i=1}^{k_{m}}\Big{|}\overline{{\mathbb{E}}}\big{[}X_{t_{i+1}^{m}}-X_{t_{i}^{m}}\big{|}{\mathcal{G}}_{T}\big{]}\Big{|}\leq\overline{{\mathbb{E}}}\bigg{[}\sum_{i=1}^{k_{m}}\big{|}X_{t_{i+1}^{m}}-X_{t_{i}^{m}}\big{|}\Big{|}{\mathcal{G}}_{T}\bigg{]},$
we have,
$\overline{{\mathbb{E}}}$ $\bigg{[}\bigg{(}\sup_{\sigma_{m}}\sum_{i=1}^{k_{m}}\big{|}X^{\mathcal{G}}_{t_{i+1}^{m}}-X^{\mathcal{G}}_{t_{i}^{m}}\big{|}\bigg{)}^{p}\bigg{]}$
$=\overline{{\mathbb{E}}}\bigg{[}\sup_{\sigma_{m}}\bigg{(}\sum_{i=1}^{k_{m}}\big{|}X^{\mathcal{G}}_{t_{i+1}^{m}}-X^{\mathcal{G}}_{t_{i}^{m}}\big{|}\bigg{)}^{p}\bigg{]}\leq\overline{{\mathbb{E}}}\bigg{[}\sup_{\sigma_{m}}\bigg{(}\overline{{\mathbb{E}}}\bigg{[}\sum_{i=1}^{k_{m}}\big{|}X_{t_{i+1}^{m}}-X_{t_{i}^{m}}\big{|}\Big{|}{\mathcal{G}}_{T}\bigg{]}\bigg{)}^{p}\bigg{]}$
$\leq\overline{{\mathbb{E}}}\bigg{[}\bigg{(}{\mathbb{E}}\bigg{[}\sup_{\sigma_{m}}\sum_{i=1}^{k_{m}}\big{|}X_{t_{i+1}^{m}}-X_{t_{i}^{m}}\big{|}\Big{|}{\mathcal{G}}_{T}\bigg{]}\bigg{)}^{p}\bigg{]}\leq\overline{{\mathbb{E}}}\bigg{[}\overline{{\mathbb{E}}}\bigg{[}\bigg{(}\sup_{\sigma_{m}}\sum_{i=1}^{k_{m}}\big{|}X_{t_{i+1}^{m}}-X_{t_{i}^{m}}\big{|}\bigg{)}^{p}\bigg{|}{\mathcal{G}}_{T}\bigg{]}\bigg{]}$
$=\overline{{\mathbb{E}}}\bigg{[}\bigg{(}\sup_{\sigma_{m}}\sum_{i=1}^{k_{m}}\big{|}X_{t_{i+1}^{m}}-X_{t_{i}^{m}}\big{|}\bigg{)}^{p}\bigg{]}=\overline{{\mathbb{E}}}\bigg{[}\bigg{(}\int_{0}^{T}\big{|}{\mathrm{d}}X_{s}\big{|}\bigg{)}^{p}\bigg{]}<\infty,$
which leads to (2a).
For (2b), due to (2a) and the decomposability (see [51, Theorem 1, Chapter 4]), it suffices to show that if $X$ is an $\mathbb{F}$}-local martingale, then $X^{\mathcal{G}}$ is a local martingale with respect to $\overline{{\mathbb{F}}}$, $\mathbb{F}$} and $\mathbb{G}$}, which thanks to (1) can be further reduced to proving that $X^{\mathcal{G}}$ is a local martingale with respect to $\mathbb{F}$} and $\mathbb{G}$}. Let $\tau_{k}$ be the stopping time $\tau_{k}:=\inf\{t\in[0,T]:[X,X]_{t}>k\}$ and consider the truncated process $X^{\tau_{k}}:=\{X_{t\wedge\tau_{k}}\}_{t\in[0,T]}$. By Burkholder-Davis-Gundy inequality,
$\overline{{\mathbb{E}}}\big{[}(X^{\tau_{k}})^{*}_{t}\big{]}^{2}\leq\overline{{\mathbb{E}}}\big{[}\big{(}(X^{\tau_{k}}\big{)}^{*}_{t})^{2}\big{]}\leq C\overline{{\mathbb{E}}}\big{[}\big{[}X^{\tau_{k}},X^{\tau_{k}}\big{]}_{t}\big{]}\leq k,$ (5.10)
for some constant $C$ and $(X^{\tau_{k}})^{*}_{t}=\sup_{0\leq s\leq t}|X_{s\wedge\tau_{k}}|.$ Let $\{\widetilde{\tau}_{\ell}\}_{\ell\in{\mathbb{N}}^{+}}$ be the sequence of the $\mathbb{F}$}-stopping times such that $\tau_{\ell}\to T$ and $X^{\tilde{\tau}_{\ell}}=\{X_{t\wedge\tilde{\tau}_{\ell}}\}_{t\in[0,T]}$ is an $\mathbb{F}$}-martingale for any $\ell\in{\mathbb{N}}^{+}$. Then $X^{\tilde{\tau}_{\ell}\wedge\tau_{k}}=\{X_{t\wedge\tilde{\tau}_{\ell}\wedge\tau_{k}}\}_{t\in[0,T]}$ is an $\mathbb{F}$}-martingale for any $\ell\in{\mathbb{N}}^{+}$. By Lemma 2.5, $|X_{t\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}|\leq(X^{\tau_{k}})^{*}_{t}$, and from (5.10),
$X_{s\wedge\tau_{k}}=\lim_{\ell\to\infty}X_{s\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}=\lim_{\ell\to\infty}\overline{{\mathbb{E}}}\big{[}X_{t\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}|{\mathcal{F}}_{s}\big{]}=\overline{{\mathbb{E}}}\Big{[}\lim_{\ell\to\infty}X_{t\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}|{\mathcal{F}}_{s}\Big{]}=\overline{{\mathbb{E}}}\big{[}X_{t\wedge\tau_{k}}|{\mathcal{F}}_{s}\big{]}.$
Combined with (5.10), we have that $X^{\tau_{k}}=\{X_{t\wedge\tau_{k}}\}_{t\in[0,T]}$ is indeed a square integrable $\mathbb{F}$}-martingale.
Now we claim that $(X^{\tau_{k}})^{\mathcal{G}}$ is a square integrable $\mathbb{G}$}-martingale. (Hence, by the H hypothesis and Remark 2.7 also an $\mathbb{F}$}-martingale). The integrability holds trivially since $X_{t}$ is integrable, and the martingale property holds since
$\overline{{\mathbb{E}}}\Big{[}X_{t\wedge\tau_{k}}^{\mathcal{G}}\big{|}{\mathcal{G}}_{s}\Big{]}$ $=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}X_{t\wedge\tau_{k}}|{\mathcal{G}}_{T}\big{]}\big{|}{\mathcal{G}}_{s}\Big{]}=\overline{{\mathbb{E}}}\big{[}X_{t\wedge\tau_{k}}|{\mathcal{G}}_{s}\big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}X_{t\wedge\tau_{k}}|{\mathcal{F}}_{s}\big{]}\big{|}{\mathcal{G}}_{s}\Big{]}$
$=\overline{{\mathbb{E}}}\big{[}X_{s\wedge\tau_{k}}|{\mathcal{G}}_{s}\big{]}=\overline{{\mathbb{E}}}\big{[}X_{s\wedge\tau_{k}}|{\mathcal{G}}_{T}\big{]}=X_{s\wedge\tau_{k}}^{\mathcal{G}}.$
This completes the proof that $X^{\mathcal{G}}$ is an $\mathbb{G}$}-local martingale and (2b).
5.3. Proof of Theorem 3.5
For (1), let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ be a sequence of partitions which tends to identity in the sense of [51, Page 64, definition] and satisfies
$\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}(X^{\mathcal{G}}_{t_{k+1}^{m}}-X^{\mathcal{G}}_{t_{k}^{m}}),\;\;\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}(X^{\prime}_{t_{k+1}^{m}}-X^{\prime}_{t_{k}^{m}}),$
where the convergence is understood as a.s. convergence. Then
$\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}$ $=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}(X^{\mathcal{G}}_{t_{k+1}^{m}}-X^{\mathcal{G}}_{t_{k}^{m}})=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}{\mathbb{E}}\Big{[}X_{t_{k+1}^{m}}-X_{t_{k}^{m}}\Big{|}{\mathcal{G}}_{T}\Big{]}$
$=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}\overline{{\mathbb{E}}}\Big{[}X^{\prime}_{t_{k+1}^{m}}-X^{\prime}_{t_{k}^{m}}\Big{|}{\mathcal{F}}\Big{]}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}{\mathbb{E}}\Big{[}Z_{t_{k}^{m}}(X^{\prime}_{t_{k+1}^{m}}-X^{\prime}_{t_{k}^{m}})\Big{|}{\mathcal{F}}\Big{]}.$
Define ${Z}^{m}$ as
$Z^{m}_{t}=Z_{0}\mathbbm{1}_{\{0\}}(t)+\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}\mathbbm{1}_{(t_{k}^{m},t_{k+1}^{m}]}(t),$
then $\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\lim_{m\to\infty}{\mathbb{E}}\big{[}\int_{0}^{t}Z_{s}^{m}{\mathrm{d}}X^{\prime}_{s}\Big{|}{\mathcal{F}}\big{]},$ and
$\sum_{k=0}^{k_{m}-1}\overline{{\mathbb{E}}}\Big{[}Z_{t_{k}^{m}}\big{(}X^{\prime}_{t_{k+1}^{m}}-X^{\prime}_{t_{k}^{m}}\big{)}\Big{]}-\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{]}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}X^{\prime}_{s}\bigg{]}.$
Recall that $X^{\prime}$ has the decomposition $X^{\prime}=M^{\prime}+V^{\prime}$, with $M^{\prime}$ a local martingale, $V^{\prime}$ an adapted finite variation process, and $\bigg{\|}\sqrt{[M^{\prime},M^{\prime}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty.$ Therefore,
$\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z^{\prime\prime}_{s-}){\mathrm{d}}X^{\prime}_{s}\bigg{|}\bigg{]}\leq\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}M^{\prime}_{s}\bigg{|}\bigg{]}+\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}V^{\prime}_{s}\bigg{|}\bigg{]}.$ (5.11)
For the first term in (5.11), applying Burkholder-Davis-Gundy inequality to get
$\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}M^{\prime}_{s}\bigg{|}\bigg{]}$ $\leq\overline{{\mathbb{E}}}\bigg{[}\sup_{0\leq s\leq t}\bigg{|}\int_{0}^{s}(Z_{u}^{m}-Z_{u-}){\mathrm{d}}M^{\prime}_{u}\bigg{|}^{p}\bigg{]}$
$\leq c_{1}\overline{{\mathbb{E}}}\bigg{[}\bigg{(}\int_{0}^{t}(Z_{u}^{m}-Z_{u-})^{2}{\mathrm{d}}[M^{\prime},M^{\prime}]_{s}\bigg{)}^{p/2}\bigg{]}$
for some constant $c_{1}$. Note that $\int_{0}^{t}(Z_{u}^{m}-Z_{u-})^{2}{\mathrm{d}}[M^{\prime},M^{\prime}]_{s}\leq 4\sup_{0\leq s\leq t}(Z_{s})^{2}[M^{\prime},M^{\prime}]_{t}$, and
$\overline{{\mathbb{E}}}\Big{[}\Big{(}\sup_{0\leq s\leq t}(Z_{s})^{2}[M^{\prime},M^{\prime}]_{t}\Big{)}^{1/2}\Big{]}\leq\bigg{(}\overline{{\mathbb{E}}}\Big{[}\sup_{0\leq s\leq t}(Z_{s})^{q}\Big{]}\overline{{\mathbb{E}}}\big{[}[M^{\prime},M^{\prime}]^{p/2}_{t}\big{]}\bigg{)}^{1/2}\leq C\sqrt{\|Z\|^{q}_{{\mathcal{H}}^{q}}\overline{{\mathbb{E}}}\big{[}[M^{\prime},M^{\prime}]_{t}^{p/2}\big{]}}<\infty.$
Since $Z$ is càdlàg, $Z^{m}_{s}\to Z_{s-},a.s.$ and $|Z^{m}_{s}-Z_{s}|\leq\sup_{0\leq u\leq s}|Z_{u}|\in S_{q}$; and Lemma 2.5 leads to
$\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}M^{\prime}_{s}\bigg{]}\to 0.$
For the second term in (5.11), we have
$\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}V^{\prime}_{s}\bigg{|}\bigg{]}\leq\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}|Z_{s}^{m}-Z_{s-}|\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\bigg{]}.$
Note that $\int_{0}^{t}|Z_{s}^{m}-Z_{s-}|\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\leq 2\sup_{0\leq s\leq t}|Z_{s}|\int_{0}^{t}|{\mathrm{d}}V_{t}|$, and similarly, we have
$\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}V^{\prime}_{s}\bigg{|}\bigg{]}\to 0,$
and
$\lim_{m\to\infty}\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}Z_{s}^{m}{\mathrm{d}}X^{\prime}_{s}-\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{|}\bigg{]}=0.$
Now define
$R^{m}:=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z^{m}_{s-}{\mathrm{d}}X^{\prime}_{s}\Big{|}{\mathcal{F}}\bigg{]},\;R:=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\Big{|}{\mathcal{F}}\bigg{]},$
for all $A\in{\mathcal{F}}$, we have
$\overline{{\mathbb{E}}}[(R-R_{m})\mathbbm{1}_{A}]=\overline{{\mathbb{E}}}\bigg{[}\bigg{(}\int_{0}^{t}Z_{s}^{m}{\mathrm{d}}X^{\prime}_{s}-\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{)}\mathbbm{1}_{A}\bigg{]}\leq\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}Z_{s}^{m}{\mathrm{d}}X^{\prime}_{s}-\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{|}\bigg{]}\to 0.$
Let $A=\{R-R_{m}>\epsilon\}$, we have
$\overline{{\mathbb{P}}}\big{(}R_{m}-R>\epsilon\big{)}\leq\frac{\overline{{\mathbb{E}}}[(R_{m}-R)\mathbbm{1}_{A}]}{\epsilon}\to 0,$
and similarly
$\overline{{\mathbb{P}}}\big{(}R_{m}-R<\epsilon\big{)}\to 0.$
Therefore, $R_{m}\to R$ in probability. Meanwhile, $R_{m}\to\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}$ a.s., hence
$\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z^{m}_{s-}{\mathrm{d}}X^{\prime}_{s}\Big{|}{\mathcal{F}}\bigg{]}.$
For (2), let us first prove that
$X^{\mathcal{G}},Y^{\mathcal{G}}]_{t}=\overline{{\mathbb{E}}}\Big{[}[X^{\prime\prime},Y^{\prime}]_{t}\Big{|}{\mathcal{G}}\Big{]}.$
By the definition of the quadratic variation,
$\big{[}X^{\mathcal{G}},Y^{\mathcal{G}}\big{]}_{t}=X^{\mathcal{G}}_{t}Y^{\mathcal{G}}_{t}-\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}-\int_{0}^{t}Y^{\mathcal{G}}_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s},$
where the last two terms are well defined because of (2b). The first term $X^{\mathcal{G}}_{t}Y^{\mathcal{G}}_{t}$ is
$X^{\mathcal{G}}_{t}Y^{\mathcal{G}}_{t}=\overline{{\mathbb{E}}}\big{[}X_{t}|{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{E}}}\big{[}Y_{t}|{\mathcal{G}}_{T}\big{]}=\overline{{\mathbb{E}}}\big{[}X^{\prime\prime}_{t}|{\mathcal{F}}\big{]}\overline{{\mathbb{E}}}\big{[}Y^{\prime}_{t}|{\mathcal{F}}\big{]}=\overline{{\mathbb{E}}}\big{[}X^{\prime\prime}_{t}Y^{\prime}_{t}|{\mathcal{F}}\big{]},$
where the last equation holds due to (3) in Proposition 5.1. For the second term $\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}$, let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ be a sequence of partitions tending to identity in the sense of [51, Page 64, definition] and satisfying
$\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}X^{\mathcal{G}}_{t_{k}^{m}}(Y^{G}_{t_{k+1}^{m}}-Y^{\mathcal{G}}_{t_{k}^{m}}),\;\;\int_{0}^{t}X^{\prime\prime}_{s-}{\mathrm{d}}Y^{\prime}_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}X^{\prime\prime}_{t_{k}^{m}}(Y^{\prime}_{t_{k+1}^{m}}-Y^{\prime}_{t_{k}^{m}}),$
where the convergence is understood as a.s. convergence. Therefore,
$\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}$ $=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}X^{\mathcal{G}}_{t_{k}^{m}}(Y^{\mathcal{G}}_{t_{k+1}^{m}}-Y^{\mathcal{G}}_{t_{k}^{m}})=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}{\mathbb{E}}\big{[}X_{t_{k}^{m}}|{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{E}}}\big{[}Y_{t_{k+1}^{m}}-Y_{t_{k}^{m}}|{\mathcal{G}}_{T}\big{]}$
$=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}{\mathbb{E}}\big{[}X^{\prime\prime}_{t_{k}^{m}}|{\mathcal{F}}\big{]}{\mathbb{E}}\big{[}Y^{\prime}_{t_{k+1}^{m}}-Y^{\prime}_{t_{k}^{m}}|{\mathcal{F}}\big{]}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}\overline{{\mathbb{E}}}\big{[}X^{\prime\prime}_{t_{k}^{m}}\big{(}Y^{\prime}_{t_{k+1}^{m}}-Y^{\prime}_{t_{k}^{m}}\big{)}|{\mathcal{F}}\big{]},$
where the last equation holds due to (3) in Proposition 5.1. Similar to the proof of (1), we conclude that
$\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}X^{\prime\prime}_{s-}{\mathrm{d}}Y^{\prime}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
and similarly
$\int_{0}^{t}Y^{\mathcal{G}}_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Y^{\prime}_{s-}{\mathrm{d}}X^{\prime\prime}_{s}\bigg{|}{\mathcal{F}}\bigg{]}.$
Therefore
$X^{\mathcal{G}},Y^{\mathcal{G}}]_{t}=\overline{{\mathbb{E}}}\Big{[}[X^{\prime\prime},Y^{\prime}]_{t}\Big{|}{\mathcal{F}}\Big{]}.$
Meanwhile, we can further assume that the partition $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ satisfies
$\int_{0}^{t}Z_{s-}{\mathrm{d}}[X^{\mathcal{G}},Y^{\mathcal{G}}]_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}\big{(}[X^{\mathcal{G}},Y^{\mathcal{G}}]_{t_{k+1}^{m}}-[X^{\mathcal{G}},Y^{\mathcal{G}}]_{t_{k}^{m}}\big{)},$
and
$\int_{0}^{t}Z_{s-}{\mathrm{d}}[X^{\prime\prime},Y^{\prime}]_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}([X^{\prime\prime},Y^{\prime}]_{t_{k+1}^{m}}-[X^{\prime\prime},Y^{\prime}]_{t_{k}^{m}}).$
Repeat the same argument as before and notice that
$\big{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}[X^{\prime\prime},Y^{\prime}]_{s}\big{|}\leq 2\sup_{0\leq s\leq t}|Z_{s}|[X^{\prime\prime},Y^{\prime}]_{t}.$
Then we have
$\overline{{\mathbb{E}}}\bigg{[}\sup_{0\leq s\leq t}|Z_{s}|[X^{\prime\prime},Y^{\prime}]_{s}\bigg{]}$ $\leq\overline{{\mathbb{E}}}[(\sup_{0\leq s\leq t}|Z_{s}|)^{r}]^{1/r}\overline{{\mathbb{E}}}\Big{[}[X^{\prime\prime},X^{\prime\prime}]_{t}^{p}\Big{]}^{1/p}\overline{{\mathbb{E}}}\Big{[}[Y^{\prime},Y^{\prime}]_{t}^{q}\Big{]}^{1/q}$
$\leq C\|Z\|_{H_{r}}\|X\|_{H_{p}}\|Y\|_{H_{q}},$
where the first inequality is by Cauchy inequality and Kunita-Watanabe inequality with $C$ a generic constant. Now repeating the arguments in (1), we achieve
$\int_{0}^{t}Z_{s-}{\mathrm{d}}\big{[}X^{\mathcal{G}},Y^{\mathcal{G}}\big{]}_{t}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}\big{[}X^{\prime\prime},Y^{\prime}\big{]}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
and similarly,
$\int_{0}^{t}Z_{s-}\big{[}X^{\mathcal{G}},Y\big{]}_{s}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z_{s-}\big{[}X^{\prime},Y\big{]}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
which completes the proof.∎
6. Application to control problems with partial observed jump-diffusions
We now apply our result (Theorem 2.11) to analyze control problems with partially observed jump-diffusions. We will derive an appropriate form of the verification theorem and recover the results in [4] as special cases.
Consider a complete filtered space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$ with a sub-filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$ satisfying the compatibility condition as in Assumption 2.6. We define as an admissible control process any $\mathbb{G}$}-progressive process $\alpha$ with values in some given Borel space $A$. The set of admissible control processes is denoted by ${\mathcal{A}}^{\mathbb{G}}$.
Given a square-integrable initial condition $\bm{\xi}\in{\mathbb{R}}^{n}$, $d$-dimensional $\mathbb{F}$}-Brownian motion $\bm{B}$, an independent Poisson random measure $N$ with a Lévy measure $\nu$, with $\widetilde{N}(\ d\bm{\theta},{\mathrm{d}}t):=N({\mathrm{d}}\bm{\theta},{\mathrm{d}}t)-\nu({\mathrm{d}}\bm{\theta}){\mathrm{d}}t$ the compensated Poisson random measure and $\alpha\in{\mathcal{A}}^{\mathbb{G}}$, the $n$-dimensional controlled state equation is defined by
${\mathrm{d}}\bm{X}_{t}^{\alpha}=\bm{b}\big{(}\bm{X}_{t}^{\alpha},\alpha_{t}\big{)}{\mathrm{d}}t+\bm{\sigma}\big{(}\bm{X}_{t}^{\alpha},\alpha_{t}\big{)}{\mathrm{d}}\bm{B}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}(\bm{X}^{\alpha}_{t-},\alpha_{t-},\bm{\theta})\widetilde{N}({\mathrm{d}}\bm{\theta},{\mathrm{d}}t),\text{ for }0\leq t\leq T,\;\bm{X}^{\alpha}_{0}=\bm{\xi}.$ (6.1)
The coefficients $\bm{b}$ and $\bm{\sigma}$ are deterministic measurable functions from ${\mathbb{R}}^{n}\times A$ into ${\mathbb{R}}^{n}$ and ${\mathbb{R}}^{n\times d}$, respectively, and the coefficient $\bm{\beta}$ are deterministic measurable functions from ${\mathbb{R}}^{n}\times A\times{\mathbb{R}}^{q}$ into ${\mathbb{R}}^{n}$ and ${\mathbb{R}}^{n\times d}$ assumed to satisfy the following standing assumptions: there exists some positive constant $C_{1}$ such that for all $\bm{x},\bm{x}^{\prime}\in{\mathbb{R}}^{n}$, $a\in A$,
$\big{|}\bm{b}(\bm{x},a)-\bm{b}(\bm{x}^{\prime},a)\big{|}+\big{|}\bm{\sigma}(\bm{x},a)-\bm{\sigma}(\bm{x}^{\prime},a)\big{|}+\int_{{\mathbb{R}}^{q}}\big{|}\bm{\beta}(\bm{x},a,\bm{\theta})-\bm{\beta}(\bm{x}^{\prime},a,\bm{\theta})\big{|}\nu({\mathrm{d}}\bm{\theta})$ $\leq C_{1}|\bm{x}-\bm{x}^{\prime}|;$
$\big{|}\bm{b}(\bm{0},a)\big{|}+\big{|}\bm{\sigma}(\bm{0},a)\big{|}+\int_{{\mathbb{R}}^{q}}\big{|}\bm{\beta}(\bm{0},a,\bm{\theta})\big{|}\nu({\mathrm{d}}\bm{\theta})$ $\leq C_{1}.$
It is clear from standard arguments that there exists a unique strong solution $\{\bm{X}_{t}^{\alpha}\}_{t\in[0,T]}$ to (6.1), which is $\mathbb{F}$}-adapted and satisfies
$\sup_{\alpha\in{\mathcal{A}}^{\mathbb{G}}}{\mathbb{E}}\bigg{[}\sup_{0\leq t\leq T}\big{|}\bm{X}_{t}^{\alpha}\big{|}^{2}\bigg{]}\leq C\Big{(}1+\textasciimacron{\mathbb{E}}\big{[}|\bm{\xi}|^{2}\big{]}\Big{)}<\infty,$
for some positive constant $C$.
The goal is to maximize, over all admissible control processes $\alpha\in{\mathcal{A}}^{\mathbb{G}}$, the following objective:
${\mathcal{J}}(\alpha):={\mathbb{E}}\bigg{[}\int_{0}^{T}f(\bm{X}_{t}^{\alpha},\alpha_{t}){\mathrm{d}}t+g(\bm{X}_{T}^{\alpha})\bigg{]}.$
Here $f:{\mathbb{R}}^{n}\times A\to{\mathbb{R}}$, and $g:{\mathbb{R}}^{n}\to{\mathbb{R}}$ are continuous functions satisfying the quadratic growth condition: there exists some positive constant $C_{2}$ such that for all $(\bm{x},a)\in{\mathbb{R}}^{n}\times A$,
$\big{|}f(\bm{x},a)\big{|}+\big{|}g(\bm{x})\big{|}\leq C_{2}\big{(}1+|\bm{x}|^{2}\big{)}.$
We can now define the objective function starting from time $t$, and for $\mu\in{\mathcal{P}}_{2}({\mathbb{R}}^{d})$ by
$J(t,\alpha,\mu):={\mathbb{E}}\bigg{[}\int_{t}^{T}f(\bm{X}_{s}^{\alpha},\alpha_{s}){\mathrm{d}}s+g(\bm{X}_{T}^{\alpha})\Big{|}\text{Law}(\bm{X}^{\alpha}_{t}|{\mathcal{G}}_{T})=\text{Law}(\bm{X}^{\alpha}_{t}|{\mathcal{G}}_{t})=\mu\bigg{]},$
and the corresponding value function by
$V(t,\mu):=\sup_{\alpha}J(t,\alpha,\mu).$
It is not hard to see that both $J$ and $V$ satisfy the quadratic growth estimate:
$|J(t,\alpha,\mu)|\leq|V(t,\mu)|\leq C\Big{(}1+\textasciimacron{\mathbb{E}}\big{[}|\bm{\xi}|^{2}\big{]}\Big{)}$
for all $\mu\in{\mathcal{P}}_{2}({\mathbb{R}}^{d}),\alpha\in{\mathcal{A}}^{\mathcal{G}}$ for some constant $C$.
Note that [4] considers the case where the information $\mathbb{G}$} is a filtration generated by a lower-dimensional Brownian motion $\bm{W}$ such that $\bm{B}=(\bm{V},\bm{W})$ and $\bm{V}$ independent of $\bm{W}$. This corresponds to the special case of (6.1) with $\bm{\beta}=0$. In this case, a form of Itô’s formula without proof is presented (see (5.7) of [4]), from which a verification theorem for the control problem with partial observation is derived. Here we will derive rigorously, as a corollary of Theorem 2.11, a generalization of (5.7) in [4] with a general form of $\mathbb{G}$} satisfying the compatibility condition.
Corollary 6.1 (Itô’s formula for flows of partial observed jump-diffusion).
[…]
$\Phi(\mu_{t})-\Phi(\mu_{0})$ (6.2)
$=$ $\;\overline{{\mathbb{E}}}\Bigg{[}\int_{0}^{t}\bigg{(}\Big{(}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}\big{)}\cdot\bm{b}(\bm{X}^{\alpha,1}_{s},\alpha_{s})+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}\big{)}:\bm{\sigma}\bm{\sigma}^{\top}\big{(}\bm{X}^{\alpha,1}_{s},\alpha_{s}\big{)}\Big{)}{\mathrm{d}}s$
$\qquad\qquad+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s},\bm{X}^{\alpha,2}_{s}\big{)}:\bm{\sigma}\big{(}\bm{X}^{\alpha,1}_{s},\alpha_{s}\big{)}\bm{\sigma}^{\top}\big{(}\bm{X}^{\alpha,2}_{s},\alpha_{s}\big{)}{\mathrm{d}}[\bm{B}^{1},\bm{B}^{2}]_{s}$
$\qquad\qquad+\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}\big{)}\cdot\bm{\sigma}(\bm{X}^{\alpha,1}_{s},\alpha_{s}){\mathrm{d}}\bm{B}^{1}_{s}\bigg{)}$
$+\sum_{0<s\leq t}\bigg{(}\frac{1}{2}\bigg{(}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s},\bm{X}^{\alpha,2}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s-},\bm{X}^{\alpha,2}_{s}\big{)}$
$\qquad\qquad\qquad-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s},\bm{X}^{\alpha,2}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s-},\bm{X}^{\alpha,2}_{s-}\big{)}\bigg{)}$
$\quad\qquad\qquad+\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s-}\big{)}\bigg{)}\mathbbm{1}_{\{\mu_{s}=\mu_{s-}\}}\bigg{|}{\mathcal{F}}\bigg{]}$
$+\sum_{0<s\leq t}\Big{(}\Phi(\mu_{s})-\Phi(\mu_{s-})\Big{)},$
for all $t\in[0,T]$, where $(\bm{X}^{\alpha,1},\bm{B}^{1},N^{1}),(\bm{X}^{\alpha,1},\bm{B}^{2},N^{2})$ satisfying for $0\leq t\leq T$,
${\mathrm{d}}\bm{X}_{t}^{\alpha,1}=\bm{b}\big{(}\bm{X}_{t}^{\alpha,1},\alpha_{t}\big{)}{\mathrm{d}}t+\bm{\sigma}\big{(}\bm{X}_{t}^{\alpha,1},\alpha_{t}\big{)}{\mathrm{d}}\bm{B}^{1}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}(\bm{X}^{\alpha,1}_{t-},\alpha_{t-},\bm{\theta})N^{1}({\mathrm{d}}\bm{\theta},{\mathrm{d}}t),$
${\mathrm{d}}\bm{X}_{t}^{\alpha,2}=\bm{b}\big{(}\bm{X}_{t}^{\alpha,2},\alpha_{t}\big{)}{\mathrm{d}}t+\bm{\sigma}\big{(}\bm{X}_{t}^{\alpha,2},\alpha_{t}\big{)}{\mathrm{d}}\bm{B}^{2}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}(\bm{X}^{\alpha,2}_{t-},\alpha_{t-},\bm{\theta})N^{2}({\mathrm{d}}\bm{\theta},{\mathrm{d}}t),$
are two conditionally independent copies of $(\bm{X}^{\alpha},\bm{B},N)$ given the sub $\sigma$-algebra ${\mathcal{G}}_{T}$ defined in an enlarged probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$, and $\overline{{\mathbb{E}}}[\cdot|{\mathcal{F}}]$ is the conditional expectation given $\mathcal{F}$} in the enlarged probability space.
HJB equations and verification theorem. In the case where $\mathbb{G}$} is the filtration generated by Brownian motion $\bm{W}$, with $\bm{B}=(\bm{V},\bm{W})$ and $\bm{V}\perp\!\!\!\perp\bm{W}$, we can rewrite the dynamic by
${\mathrm{d}}\bm{X}_{t}^{\alpha}=\bm{b}\big{(}\bm{X}_{t}^{\alpha},\alpha_{t}\big{)}{\mathrm{d}}t+\bm{\sigma}^{\bm{V}}\big{(}\bm{X}_{t}^{\alpha},\alpha_{t}\big{)}{\mathrm{d}}\bm{V}_{t}+\bm{\sigma}^{\bm{W}}\big{(}\bm{X}_{t}^{\alpha},\alpha_{t}\big{)}{\mathrm{d}}\bm{W}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}(\bm{X}^{\alpha}_{t-},\alpha_{t-},\bm{\theta})\widetilde{N}({\mathrm{d}}\bm{\theta},{\mathrm{d}}t),$
and the conditionally independent copies
${\mathrm{d}}\bm{X}_{t}^{\alpha,1}=\bm{b}\big{(}\bm{X}_{t}^{\alpha,1},\alpha_{t}\big{)}{\mathrm{d}}t+\bm{\sigma}^{\bm{V}}\big{(}\bm{X}_{t}^{\alpha,1},\alpha_{t}\big{)}{\mathrm{d}}\bm{V}^{1}_{t}+\bm{\sigma}^{\bm{W}}\big{(}\bm{X}_{t}^{\alpha,1},\alpha_{t}\big{)}{\mathrm{d}}\bm{W}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}(\bm{X}^{\alpha,1}_{t-},\alpha_{t-},\bm{\theta})\widetilde{N}^{1}({\mathrm{d}}\bm{\theta},{\mathrm{d}}t),$
${\mathrm{d}}\bm{X}_{t}^{\alpha,2}=\bm{b}\big{(}\bm{X}_{t}^{\alpha,2},\alpha_{t}\big{)}{\mathrm{d}}t+\bm{\sigma}^{\bm{V}}\big{(}\bm{X}_{t}^{\alpha,2},\alpha_{t}\big{)}{\mathrm{d}}\bm{V}^{2}_{t}+\bm{\sigma}^{\bm{W}}\big{(}\bm{X}_{t}^{\alpha,2},\alpha_{t}\big{)}{\mathrm{d}}\bm{W}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}(\bm{X}^{\alpha,2}_{t-},\alpha_{t-},\bm{\theta})\widetilde{N}^{2}({\mathrm{d}}\bm{\theta},{\mathrm{d}}t).$
Note that under this setting, at time $s$, $\bm{X}^{\alpha,1}_{s-}$ and $\bm{X}^{\alpha,2}_{s-}$ has a jump $\bm{\beta}(\bm{X}^{\alpha,1}_{s-},\alpha_{s-},\theta^{1}),\bm{\beta}(\bm{X}^{\alpha,2}_{s-},\alpha_{s-},\theta^{2})$ with Poisson random measure $N^{1}({\mathrm{d}}\theta^{1},{\mathrm{d}}s)\times N^{2}({\mathrm{d}}\theta^{2},{\mathrm{d}}s)$, we can simplify the Itô’s formula by
$\Phi(\mu_{t})-\Phi(\mu_{0})$ (6.3)
$=$ $\;\int_{0}^{t}\overline{{\mathbb{E}}}\Big{[}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}\big{)}\cdot\bm{b}(\bm{X}^{\alpha,1}_{s},\alpha_{s})+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}\big{)}:\bm{\sigma}\bm{\sigma}^{\top}\big{(}\bm{X}^{\alpha,1}_{s},\alpha_{s}\big{)}$
$\qquad\qquad+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s},\bm{X}^{\alpha,2}_{s}\big{)}:\bm{\sigma}^{\bm{W}}\big{(}\bm{X}^{\alpha,1}_{s},\alpha_{s}\big{)}{\bm{\sigma}^{\bm{W}}}^{\top}\big{(}\bm{X}^{\alpha,2}_{s},\alpha_{s}\big{)}\Big{|}{\mathcal{F}}\Big{]}{\mathrm{d}}s$
$\quad+\frac{1}{2}\int_{0}^{t}\int_{{\mathbb{R}}^{2q}}\overline{{\mathbb{E}}}\Big{[}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}+\bm{\beta}(\bm{X}^{\alpha,1}_{s},\alpha_{s},\bm{\theta}^{1}),\bm{X}^{\alpha,2}_{s}+\bm{\beta}(\bm{X}^{\alpha,2}_{s},\alpha_{s},\bm{\theta}^{2})\big{)}$
$-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s},\bm{X}^{\alpha,2}_{s}+\bm{\beta}(\bm{X}^{\alpha,2}_{s},\alpha_{s},\bm{\theta}^{2})\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}+\bm{\beta}(\bm{X}^{\alpha,1}_{s},\alpha_{s},\bm{\theta}^{1}),\bm{X}^{\alpha,2}_{s}\big{)}$
$\qquad\qquad\qquad\qquad+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s},\bm{X}^{\alpha,2}_{s}\big{)}\Big{|}{\mathcal{F}}\Big{]}\nu({\mathrm{d}}\bm{\theta}^{1})\nu({\mathrm{d}}\bm{\theta}^{2}){\mathrm{d}}s$
$\quad\qquad+\int_{0}^{t}\int_{{\mathbb{R}}^{q}}\overline{{\mathbb{E}}}\Big{[}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s}+\bm{\beta}(\bm{X}^{\alpha,1}_{s},\alpha_{s},\bm{\theta})\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s}\big{)}$
$\qquad\qquad\qquad\qquad\quad-\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s}\big{)}\cdot\bm{\beta}(\bm{X}^{\alpha,1}_{s},\alpha_{s},\bm{\theta})\Big{|}{\mathcal{F}}\Big{]}\nu({\mathrm{d}}\bm{\theta}){\mathrm{d}}s$
$+\int_{0}^{t}\overline{{\mathbb{E}}}\Big{[}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}\big{)}\cdot\bm{\sigma}^{\bm{W}}(\bm{X}^{\alpha,1}_{s},\alpha_{s})\Big{|}{\mathcal{F}}\Big{]}{\mathrm{d}}\bm{W}_{s},$
Proceeding by the same argument as in [20], we have the dynamic programming principle for the control problem with partial observation. In particular, for all $0\leq t<s\leq T$,
$V(t,\mu)=\inf_{\alpha\in{\mathcal{A}}^{G}}{\mathbb{E}}\bigg{[}\int_{t}^{s}f(\bm{X}^{\alpha}_{u},\alpha_{u}){\mathrm{d}}u+V\big{(}s,\text{Law}(\bm{X}_{s}^{\alpha}|{\mathcal{G}}_{T})\big{)}\bigg{|}\text{Law}(\bm{X}_{t}^{\alpha}|{\mathcal{G}}_{T})=\mu\bigg{]}.$
We have now the ingredients for deriving the Bellman equation. For $\Phi\in C^{2,2}({\mathcal{P}}_{2}({\mathbb{R}}^{n}))$, $\mu\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$ and $a\in A$, the operator $\mathcal{L}^{\mu,a}$ maps from $C^{2}({\mathcal{P}}_{2}({\mathbb{R}}^{n}))$ to a function ${\mathbb{R}}^{n}\to{\mathbb{R}}$ defined by
$\big{(}\mathcal{L}^{\mu,a}\Phi\big{)}(\bm{x}):=$ $\;\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu,\bm{x}\big{)}\cdot\bm{b}(\bm{x},a)+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu,\bm{x}\big{)}:\bm{\sigma}\bm{\sigma}^{\top}\big{(}\bm{x},a\big{)}$
$+\int_{{\mathbb{R}}^{q}}\bigg{(}\frac{\delta\Phi}{\delta\mu}\big{(}\mu,\bm{x}+\bm{\beta}(\bm{x},a,\bm{\theta})\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu,\bm{x}\big{)}-\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu,\bm{x}\big{)}\cdot\bm{\beta}(\bm{x},a,\bm{\theta})\bigg{)}\nu({\mathrm{d}}\bm{\theta}).$
Furthermore, for $\Phi\in C^{2,2}({\mathcal{P}}_{2}({\mathbb{R}}^{n}))$, $\mu\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$, and $a\in A$, the operator $\mathcal{M}^{\mu,a}$ maps from $C^{2,2}({\mathcal{P}}_{2}({\mathbb{R}}^{n}))$ to a function ${\mathbb{R}}^{2n}\to{\mathbb{R}}$ defined by
$\big{(}\mathcal{M}^{\mu,a}\Phi\big{)}(\bm{x},\bm{x}^{\prime}):=$ $\;\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu,\bm{x},\bm{x}^{\prime}\big{)}:\bm{\sigma}^{\bm{W}}\big{(}\bm{x},a\big{)}{\bm{\sigma}^{\bm{W}}}^{\top}\big{(}\bm{x}^{\prime},a\big{)}\Big{)}$
$+\int_{{\mathbb{R}}^{2q}}\bigg{(}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu,\bm{x}+\bm{\beta}(\bm{x},a,\bm{\theta}^{1}),\bm{x}^{\prime}+\bm{\beta}(\bm{x}^{\prime},a,\bm{\theta}^{2})\big{)}$
$-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu,\bm{x},\bm{x}^{\prime}+\bm{\beta}(\bm{x}^{\prime},a,\bm{\theta}^{2})\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu,\bm{x}+\bm{\beta}(\bm{x},a,\bm{\theta}^{1}),\bm{x}^{\prime}\big{)}$
$\qquad\qquad+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu,\bm{x},\bm{x}^{\prime}\big{)}\bigg{)}\nu({\mathrm{d}}\bm{\theta}^{1})\nu({\mathrm{d}}\bm{\theta}^{2}).$
Now, the Bellman equation takes the following form:
$\begin{cases}\partial_{t}v(t,\mu)+\sup_{a\in A}\Big{(}\big{\langle}\mu,f(\cdot,a)+\mathcal{L}^{\mu,a}v(t,\mu)(\cdot)\big{\rangle}+\big{\langle}\mu\otimes\mu,\mathcal{M}^{\mu,a}v(t,\mu)(\cdot,\cdot)\big{\rangle}\Big{)}=0,\\ v(T,\mu)=\big{\langle}\mu,g\big{\rangle}.\end{cases}$ (6.4)
The Bellman equation (6.4) is validated by the following verification theorem, which can be proved by following the similar argument as in [4], with the Itô’s formula (6.3).
Theorem 6.2.
Let $v$ be a real-valued function in $C^{1,2}([0,T]\times{\mathcal{P}}_{2}({\mathbb{R}}^{n}))$, i.e., $v$ is continuous on $0,T]\times{\mathcal{P}}_{2}({\mathbb{R}}^{n})$, $v(t,\cdot)\in C^{2}({\mathcal{P}}_{2}({\mathbb{R}}^{n}))$ for all $t\in[0,T]$, and $v(\cdot,\mu)\in C^{1}([0,T))$ for all $\mu\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$. Suppose that $v$ is a solution to the Bellman equation (6.4), and there exists for all $(t,\mu)\in[0,T)\times{\mathcal{P}}_{2}({\mathbb{R}}^{n})$ an element $\hat{a}(t,\mu)$ attaining the supremum in (6.4) such that for a given square integrable initial condition, the stochastic McKean-Vlasov SDE
${\mathrm{d}}\hat{\bm{X}}_{t}=$ $\bm{b}\Big{(}\hat{\bm{X}}_{t},\hat{a}\big{(}t,\text{Law}(\hat{\bm{X}_{t}}|\bm{W})\big{)}\Big{)}{\mathrm{d}}t+\bm{\sigma}^{\bm{V}}\Big{(}\hat{\bm{X}}_{t},\hat{a}\big{(}t,\text{Law}(\hat{\bm{X}_{t}}|\bm{W})\big{)}\Big{)}{\mathrm{d}}\bm{V}_{t}$
$+\bm{\sigma}^{\bm{W}}\Big{(}\hat{\bm{X}}_{t},\hat{a}\big{(}t,\text{Law}(\hat{\bm{X}_{t}}|\bm{W})\big{)}\Big{)}{\mathrm{d}}\bm{W}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}\Big{(}\hat{\bm{X}}_{t},\hat{a}\big{(}t,\text{Law}(\hat{\bm{X}_{t}}|\bm{W})\big{)},\bm{\theta}\Big{)}N({\mathrm{d}}\bm{\theta},{\mathrm{d}}t)$
admits a unique solution $\hat{\bm{X}}$. Then, $v=V$ and the feedback control in ${\mathcal{A}}^{\mathcal{G}}$ defined by
$\alpha^{\star}_{s}:=\hat{a}\big{(}s,\text{Law}(\hat{\bm{X}_{s}}|\bm{W})\big{)},\text{ for }t\leq s\leq T,$
is an optimal control for $V(t,\mu)$. That is, $V(t,\mu)=J(t,\alpha^{\star},\mu)$.
Note that this generalizes the verification results in [4] to the case of jump-diffusions.